### v1

# Early-Exit Graph Neural Networks

Abstract
Early-exit mechanisms allow deep neural networks to halt inference as soon as classification confidence is high enough, adaptively trading depth for confidence, and thereby cutting latency and energy on easy inputs while retaining full-depth accuracy for harder ones. Similarly, adding early exit mechanisms to Graph Neural Networks (GNNs), the go-to models for graph-structured data, allows for dynamic trading depth for confidence on simple graphs while maintaining full-depth accuracy on harder and more complex graphs to capture intricate relationships. Although early exits have proven effective across various deep learning domains, their potential within GNNs in scenarios that require deep architectures while resisting over-smoothing and over-squashing remains largely unexplored. We unlock that potential by first introducing Symmetric-Anti-Symmetric Graph Neural Networks (SAS-GNN), whose symmetry-based inductive biases mitigate these issues and yield stable intermediate representations that can be useful to allow early exiting in GNNs. Building on this backbone, we present Early-Exit Graph Neural Networks (EEGNNs), which append confidence-aware exit heads that allow on-the-fly termination of propagation based on each node or the entire graph. Experiments show that EEGNNs preserve robust performance as depth grows and deliver competitive accuracy on heterophilic and long-range benchmarks, matching attention-based and asynchronous message-passing models while substantially reducing computation and latency. We plan to release the code to reproduce our experiments.
Department of Computer Science, Control and Management Engineering,
Sapienza University of Rome, Rome, Italy
Institute of Information Science and Technologies "Alessandro Faedo" - ISTI-CNR, Pisa, Italy
Corresponding author: difrancesco@diag.uniroma1.it
§Information Engineering Department, University of Pisa, Pisa, Italy
1 Introduction
Deep learning models are increasingly deployed in latency and energy-constrained settings (e.g., mobile AR, autonomous drones, real-time recommendation). Graph Neural Networks (GNNs) inherit these constraints because their message-passing depth directly translates into runtime and energy costs. In such scenarios, adapting computational effort to input difficulty is critical for both efficiency and sustainability. Graph Neural Networks (GNNs) have emerged as a powerful class of deep learning models designed to process graph-structured data. Graphs are a natural way to represent information across various domains, such as text, audio, images, knowledge representation, and social networks (Ul Qumar et al., 2023; van den Berg et al., 2017; Hamaguchi et al., 2018). The ability of GNNs to model these complex structures has allowed them to excel in tasks such as node and graph classification (Xu et al., 2019), as well as edge prediction (Kumar et al., 2020). Most GNNs follow the message-passing paradigm (Gilmer et al., 2017), and they are referred to as Message-Passing Neural Networks (MPNNs). In this paradigm, each layer updates the representation of a node by aggregating feature information from its immediate neighbors through graph convolutions, attention mechanisms, or learned neural functions (Veličković, 2023). Consequently, the number of layers of an MPNN dictates the length of paths from which information from nodes can propagate. Intuitively, increasing the number of layers should allow for better integration of long-range information, enabling messages to traverse farther across the graph and enriching each node’s representation with broader structural context. In practice, however, depth is a double-edged sword: too many layers trigger over-smoothing, in which node embeddings become indistinguishable, and over-squashing, where information from distant nodes is aggressively compressed (NT and Maehara, 2019; Topping et al., 2022); too few layers lead to under-reaching, where messages cannot cover the task-specific problem radius (Alon and Yahav, 2021). Because the problem radius cannot be known a priori, the layer count becomes a delicate hyperparameter that should be set to the shallowest depth still able to span the necessary receptive field while keeping model size and training cost as low as possible. However, even if we were able to set the number of layers to its optimal value, it is known that Message-Passing Neural Networks remain no more expressive than the 1-Weisfeiler–Lehman test (Weisfeiler and Lehman, 1968; Xu et al., 2019), limiting their ability to distinguish certain graph structures regardless of depth.
Our idea. Rather than selecting a single “best” depth, we let the network decide on-the-fly: we endow GNNs with early exits so that each node (or the whole graph) can halt message passing as soon as its prediction is confident. We let each node (or the whole graph) decide how far messages need to travel by attaching confidence-aware exit heads to a depth-stable backbone. However, early exits are reliable only if intermediate representations remain both informative and parameter-efficient; achieving this requires a new backbone. To that end, we design Symmetric–Anti-Symmetric GNNs (SAS-GNNs), whose weight-shared, ODE-inspired message passing produces stable embeddings and constant memory usage—making early exit practical at scale.
Contributions.
  1. 1.
    EEGNN. First end-to-end early-exit GNN: each node or the whole graph halts message passing when confident via Gumbel–Softmax heads, reducing inference time and removing depth tuning.
  2. 2.
    SAS-GNN backbone. A weight-shared, symmetry/anti-symmetry MPNN (Neural ODE-inspired) that offers stable mid-layer states for safe early exits with constant memory.
  3. 3.
    Theory. We prove SAS-GNN preserves node-wise information and creates controllable attraction/repulsion, countering over-smoothing and over-squashing—thus supporting reliable exits.
  4. 4.
    Results. On heterophilic benchmarks and LRGB, EEGNN/SAS-GNN match or beat attention- and asyncronous-based MPNNs with far fewer parameters, no normalization/dropout, and lower latency.
2 Related Work
Early-Exit for GNNs. Our approach shares similarities with Spinelli et al. (2021), which introduces an early-exit mechanism for GNNs, allowing nodes to halt their updates during message passing. While their method considers over-smoothing (OST), no solution for over-squashing (OSQ) was proposed. Additionally, their design requires auxiliary loss terms to enable differentiable node-level exit decisions. Other works have explored early exit mechanisms in GNNs, but these do not investigate theoretically either OST or OSQ (Xiao et al., 2021; Han et al., 2024).
In contrast, we are the first to propose early-exit not to bypass OST and OSQ. For instance, our method enables nodes to remain active in the network as long as necessary for feature extraction, naturally addressing both OST and OSQ through its architecture.
Then, our early-exit mechanism is fully differentiable and trained end-to-end using only the task loss, without the need for additional regularization or supervision signals.
While all of the above works considering early-exit have focused solely on node-level tasks, we are the first to include an extension for graph classification/regression.
Asynchronous / Transformer GNNs. Graph Transformers (GTs) (Shi et al., 2021) use self-attention to enable unrestricted node interactions, avoiding reliance on graph topology or increased depth to model long-range dependencies. However, this comes at a cost of quadratic time complexity in the number of nodes, limiting scalability. While GTs seem promising, Cai et al. (2023) shows they are not inherently more expressive than MPNNs with a virtual node, which also allows global communication. Then, in practice, GTs do not consistently outperform classical MPNNs like GCN (Kipf and Welling, 2017) on the LRGB benchmark (Tönshoff et al., 2023), casting doubt on their effectiveness for long-range tasks. Moreover, both GTs and MPNNs often rely on normalization and dropout (Luo et al., 2024), which can obscure the model’s internal behavior. Asynchronous MPNNs (Finkelshtein et al., 2024; Errica et al., 2023) dynamically adapt the graph topology at each layer to track specific paths between distant nodes and improve long-range communication, though at the cost of increased architectural complexity. They are opposed to synchronous MPNNs that operate on a fixed topology. Co-GNN (Finkelshtein et al., 2024) introduces node-level interaction modes (e.g., isolate, broadcast, listen) to address OST and OSQ. Inspired by this, we also allow discrete node decisions—but in our model, we choose whether to exit or continue, preserving the original topology and ensuring efficiency. Unlike Co-GNNs, which require manual depth tuning and are costlier at inference, our method is lightweight and scalable. AMP (Errica et al., 2023) learns depth and a message filtering strategy to address OST, OSQ, and under-reaching, but applies a fixed depth at inference. In contrast, our model adapts depth at both training and test time while using a fixed topology.
Neural ODEs methods. A smaller subset of MPNNs belongs to the class of Graph Neural Ordinary Differential Equations (ODEs) (Poli et al., 2021). These approaches aim to design more principled architectures by avoiding components like dropout and normalization layers, whether not supported by theory. They also tend to be memory-efficient, as weight matrices are typically shared across layers. Although models in this family have been proposed to address OST or OSQ individually (Di Giovanni et al., 2023; Gravina et al., 2023), to the best of our knowledge, no method has tackled both phenomena simultaneously within this framework, neither have they been used to design early-exit GNN architectures.
An extended version of the related works, covering GNNs in general and Early-Exit Neural Networks, is in Appendix G.
3 Methodology
We use bold fonts for both matrices and vectors, with uppercase letters representing matrices and lowercase letters representing vectors (e.g., $\mathbf{M}$, $\mathbf{v}$). Scalars are denoted by italic letters (e.g., s). Notation. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ be an undirected graph, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\mathbf{X}\in\mathbb{R}^{n\times m}$ is the instance matrix containing vector representations of features for each node. The $u$-$th$ row of the instance matrix is represented by $\mathbf{x}_{u}\in\mathbb{R}^{m}$. The number of nodes in $\mathcal{G}$ is denoted by $|\mathcal{V}|=n$, and the number of edges by $|\mathcal{E}|$. The symbol $\Gamma(u)$ represents the neighborhood of node $u$, and $|\Gamma(u)|$ denotes its degree. The diagonal matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ is such that the $u$-$th$ element on the diagonal is equal to $|\Gamma(u)|$. The set of edges $\mathcal{E}$ can also be expressed as the adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$, where $A_{uv}=1$ if $(u,v)\in\mathcal{E}$, and $A_{uv}=0$ otherwise. Graph Neural Networks. Most GNNs utilize the message-passing formalism to extract features from graph-structured datasets (Gilmer et al., 2017). In this formalism, the instance matrix $\mathbf{X}$ is iteratively transformed across layers during the forward pass of the GNN. We define the initial hidden representation as $\mathbf{H}^{0}=\mathbf{X}$, or as $\mathbf{H}^{0}=f(\mathbf{X})$ when using a learnable projection function $f$, such as a multi-layer perceptron (MLP). At each layer $l$, the model computes an intermediate representation matrix $\mathbf{H}^{l}$ for $0\leq l\leq L$, where the feature vector of node $u$ is denoted by $\mathbf{h}_{u}^{l}\in\mathbb{R}^{m^{\prime}}$, with $m^{\prime}$ representing the hidden dimension. This process is conditioned at each layer by $\mathbf{A}$ up to layer $L$, yielding the final node representations $\mathbf{H}^{L}\in\mathbb{R}^{n\times m^{\prime}}$, which we denote as $\mathbf{Z}$. Once we have $\mathbf{Z}$, it is used for downstream tasks such as node or graph classification, typically through a specific readout function, followed by an MLP $g(\cdot)$. The former is given by $\hat{y}_{v}=g(\mathbf{z}_{v})$ and the latter by $\hat{y}=g(\text{Pool}(\mathbf{Z}))$, where Pool (e.g.,  mean or max) is any permutation-invariant aggregator. When considering multiple graphs, we denote each graph $i$ in a dataset $\mathcal{D}=\{\mathcal{G}_{i}:i\in\mathcal{D}\}$ where $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i},\mathbf{X}_{i})$, with associated adjacency matrix $\mathbf{A}_{i}$. Similarly, we define the hidden representation matrix at layer $l$ for graph $i$ as $\mathbf{H}^{l}_{i}$. When possible, we describe the approach for a single graph. We define a task as transductive when it is performed on a single graph, while the task is inductive when datasets contain more than one graph, namely those in the LRGB.
Our approach. A naïve scheme would let every node draw an $\arg\max$ over the two actions—exit or continue—at each layer and stop if it chooses the first. Unfortunately, the hard $\arg\max$ is non-differentiable, so the exit policy cannot be learned with standard back-propagation.
To retain end-to-end differentiability, we therefore need a soft, trainable substitute for the discrete decision. This requirement leads to two concrete design goals:
  1. O1.
    Stable backbone. Design the message-passing backbone so that hidden node features stay stable and distinct as layers accumulate, i.e., they neither blow up nor collapse into identical vectors. With useful information preserved at every depth, any layer can serve as a trustworthy early-exit point.
  2. O2.
    Contextual exit policy. Equip each layer with a differentiable confidence head (implemented here with the Gumbel–Softmax trick) that decides, on the fly, whether a node or the whole graph has gathered enough evidence to stop.
Because early-exit heads can only act on reliable hidden states, O1 is a prerequisite for O2.
Symmetric-Anti-Symmetric Graph Neural Network (O1).
To accomplish O1, we build upon two message-passing schemes proposed in recent years. One is the Anti-Symmetric Deep Graph Network (A-DGN) (Gravina et al., 2023), and the other is the Gradient Flow Framework (GRAFF) (Di Giovanni et al., 2023); full details are provided in Appendix B. The former was introduced to prevent OSQ by leveraging antisymmetric learnable matrices. This results in a GNN that behaves as a stable and non-dissipative dynamical system—i.e., the real parts of each Jacobian’s eigenvalue are zero, namely non-positive (stability), and non-negative (non-dissipativeness). Concretely, this ensures that the gradient $\partial\mathbf{h}_{i}^{l}/\partial\mathbf{h}_{i}^{0}$ for each node $i$ remains well-conditioned across layers, neither vanishing nor exploding. This enables long-range dependencies to be preserved during propagation, which is essential for tasks where the problem radius is large. Further discussion is available in Appendix B. The latter, instead, was introduced to deal with node classification in heterophilic graphs [^1] A graph is heterophilic when adjacent nodes tend to share different class labels.. Since OST causes nodes to converge to the same representation, its negative effect is more enhanced when learning in heterophilic graphs, where adjacent nodes with different classes should have a distinguished representation. GRAFF takes advantage of symmetric learnable matrices, which provably enable attraction and repulsion edge-wise to prevent adjacent nodes from becoming similar in the limit of many layers. Since these approaches rely uniquely on the parameter design level, our theoretical contribution is to devise a unique architecture that takes advantage of both methods. Specifically, we designed a message-passing scheme that accounts for OST through symmetric weight matrices and OSQ through the antisymmetric counterpart.
We name such architecture Symmetric-Anti-Symmetric Graph Neural Network (SAS-GNN). Its message-passing rule is
$\dot{\mathbf{H}}^{t}=\sigma_{1}(-\sigma_{2}(\mathbf{H}^{t}\mathbf{\Omega}_{as})+\mathbf{\bar{A}}\mathbf{H}^{t}\mathbf{W}_{s}),$ (1)
where we have, without loss of generalization, $\mathbf{H}^{t}\equiv\mathbf{H}^{l}$, $\mathbf{\bar{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, that is the normalized adjacency matrix. We refer to $\sigma_{1},\sigma_{2}$ as the non-linear activation functions, and finally $\mathbf{\Omega}_{as}$, $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are the antisymmetric and symmetric trainable weight matrices. We formalised Equation (1) as an ODE since we build upon the Graph Neural ODE framework. We can integrate it via the Euler discretization as:
$\mathbf{H}^{t+\tau}=\mathbf{H}^{t}+\tau\sigma_{1}(-\sigma_{2}(\mathbf{H}^{t}\mathbf{\Omega}_{as})+\mathbf{\bar{A}}\mathbf{H}^{t}_{i}\mathbf{W}_{s})$ (2)
Here, $\tau$ is the integration step. Weight matrices are shared across layers, as specified in GRAFF. A-DGN performs similarly with or without weight sharing (Gravina et al., 2023), but we adopt it for space efficiency when scaling to many layers ($t\rightarrow\infty$). This design is supported by the following theorems.
Theorem 3.1.
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). Assuming that $\bar{\mathbf{A}}$ does not contain self-loops, and the derivative of $\sigma_{1}$ is bounded, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative.
Proof sketch. This result follows from a standard stability analysis of the ODE in Equation (1). The antisymmetric term $-\sigma_{2}(\mathbf{H}^{t}\boldsymbol{\Omega}_{\mathrm{as}})$ contributes Jacobian eigenvalues with purely imaginary components, while the symmetric term involving $\mathbf{W}_{s}$ does not increase their real parts.
Theorem 3.2.
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). Assuming that $\sigma_{1}$, and $\sigma_{2}$ are defined s.t. $\forall x\in\mathbb{R}$, $\sigma_{1}(x),\sigma_{2}(x)\geq 0$. The evolution of $\mathbf{H}^{t}$ minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
The functional that we refer to is the following
$E_{\theta}(\mathbf{H}^{t})=-\sum_{i,j}\frac{1}{\sqrt{d_{i}\cdot d_{j}}}\langle\mathbf{h}_{i}^{t},\mathbf{W}_{s}\mathbf{h}_{j}^{t}\rangle.$ (3)
Proof sketch. We show that Equation (1) monotonically decreases the energy functional in Equation (3). The symmetry of $\mathbf{W}_{s}$ allows its decomposition into eigencomponents, revealing how the dynamics induce attraction or repulsion between nodes. The edge-wise interpretation is detailed in Equation 28 (Appendix).
Corollary 3.3.
Let $\sigma_{1}(x)=\text{ReLU}(\tanh(x))$ and $\sigma_{2}(x)=\text{ReLU}(x)$, both of which are non-negative functions. Given that the derivative of $\sigma_{1}(x)$ is bounded, the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative. Furthermore, this evolution minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction and repulsion among adjacent nodes.
Proof sketch. The proof, follows directly from Theorems 3.1 and 3.2, as ReLU+TanH and ReLU satisfy the required boundedness and non-negativity conditions. We adopt this activation pair in all experiments, and show in Appendices F.3 and F.6 that it impacts the performance.
Since the LRGB made available edge features for each graph, we also propose a SAS-GNN version that encompasses their use $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ as follows.
$\dot{\mathbf{H}}^{t}=\sigma_{1}(-\sigma_{2}(\mathbf{H}^{t}\mathbf{\Omega}_{as})+f_{e}(\mathbf{E})+\mathbf{\bar{A}}\mathbf{H}^{t}\mathbf{W}_{s})$ (4)
Hu et al. (2020) implements $f_{e}(\mathbf{E})=\mathbf{B}\mathbf{E}\mathbf{W}_{e}$, where $\mathbf{B}\in\mathbb{R}^{n\times|\mathcal{E}|}$ is the node-edges incidence matrix, s.t. $B_{i,(u,v)}=1$ if $i=u\vee i=v$, otherwise $B_{i,(u,v)}=0$, and $\mathbf{W}_{e}\in\mathbb{R}^{d\times m^{\prime}}$ is a learnable weight matrix. In this work, we propose $f_{e}(\mathbf{E})\equiv-ReLU(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$. Since $\mathbf{E}$ is not dependent on $t$ or the node features, we can state the following theorem.
Theorem 3.4.
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to equation (4). If $\sigma_{1}(x)=\text{ReLU}(\text{tanh}(x))$, $\sigma_{2}(x)=\text{ReLU}(x)$, and $f_{e}(\mathbf{E})\equiv-ReLU(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative and minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
This theorem can be proved equivalently to the previous ones. Theorems’ proofs are included in Appendix C.
Algorithm 1 Neural Adaptive-Step Early-Exit GNNs for Node Classification
1:Initialize $\mathbf{H}^{0}$, $L$, $\mathbf{\bar{A}}$, $f_{e}$, $f_{c}$, $f_{\nu}$, $\mathbf{W}_{s}$, $\mathbf{\Omega}_{as}$, $\sigma_{1}$, $\sigma_{2}$, $\nu_{0}$, $\mathbf{Z}=\mathbf{0}_{n\times d}$, $exit\_list=\{\}$
2:for $l=0$ to $L$ do
3:     $\mathbf{C}^{l}\leftarrow f_{c}(\mathbf{H}^{l})$; $\bm{\nu}^{l}\leftarrow f_{\nu}(\mathbf{H}^{l},\nu_{0})$
4:     $\mathbf{c}^{l}\leftarrow gumbel\_softmax(\mathbf{C}^{l},\bm{\nu}^{l})$
5:     $\bm{\tau}^{l}\leftarrow\mathbf{c}^{l}(\texttt{0})$
6:     $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}+\bm{\tau}^{l}\sigma_{1}(-\sigma_{2}(\mathbf{H}^{l}\mathbf{\Omega}_{as})+f_{e}(\mathbf{E})+\mathbf{\bar{A}}\mathbf{H}^{l}\mathbf{W}_{s})$
7:     for $i=0$ to $n$ do
8:         if $argmax\{\mathbf{c}^{l}\}=\texttt{1}\wedge i\notin exit\_list$ then
9:              $\mathbf{Z}_{i}\leftarrow\mathbf{h}_{i}^{l}$; $exit\_list.add(i)$
10:for $i=0$ to $n$ do
11:     if $i\notin exit\_list$ then
12:         $\mathbf{Z}_{i}\leftarrow\mathbf{h}_{i}^{L}$;$exit\_list.add(i)$
13:return $\mathbf{Z}$
Gumbel Softmax Early-Exit Mechanism (O2).
Having addressed O1, we now turn to O2: implementing a contextual early-exit mechanism. We first focus on node classification; the extension to graph classification is presented in Appendix D. Recently, a very similar problem was faced by Finkelshtein et al. (2024), where each node had to take actions based on the context. Their strategy employed the straight-through Gumbel-Softmax estimator (Jang et al., 2017; Maddison et al., 2017), which derives a continuous approximation of discrete action sampling. Taking inspiration from Finkelshtein et al. (2024), we propose something similar but with a reduced action space. Let $\Omega$ be the action space, we have $|\Omega|=2$, where $\Omega=\{\texttt{0},\texttt{1}\}$, with 0 is the non-exiting, and 1 is the exiting action. In our work, we define $\mathbf{C}\in\mathbb{R}^{|\Omega|}$ as a confidence vector representing action probabilities. For example, $\mathbf{C}(\texttt{1})$ gives the probability of exiting at a specific network state. The Gumbel-Softmax estimator approximates the categorical distribution $\mathbf{C}$ using a Gumbel-distributed vector $\mathbf{g}\in\mathbb{R}^{|\Omega|}$, where each component $g(a)\sim\text{GUMBEL}(0,1)$ for $a\in\{\texttt{0},\texttt{1}\}$. Given $\mathbf{C}$ and a temperature parameter $\nu$, the Gumbel-Softmax score for node $i$ is computed as:
$\mathbf{c}_{i}(\mathbf{C}_{i};\nu_{i})=\nicefrac{{\exp\left(\frac{\log(\mathbf{C}_{i})+\mathbf{g}_{i}}{\nu_{i}}\right)}}{{\sum_{a\in\{\texttt{0},\texttt{1}\}}\exp\left(\frac{\log(C_{i}(a))+g_{i}(a)}{\nu_{i}}\right)}},$
we observe that as $\nu_{i}\to 0$, this expression approaches a one-hot encoding. We compute $\mathbf{C}_{i}$ and $\nu_{i}$ from the current context $\mathbf{H}^{t}$ in matrix form as $\mathbf{C}^{t}\in\mathbb{R}^{n\times|\Omega|}$ and $\bm{\nu}^{t}\in\mathbb{R}^{n\times 1}$, via $f_{c}$ and $f_{\nu}$ which are neural networks with fixed depth $L_{f}$ and hidden dimension $m_{f}$. For efficiency purposes, $f_{c}$ and $f_{\nu}$ are shared across layers, even though their outputs $\mathbf{C}^{t}$ and $\bm{\nu}^{t}$ depend on time $t$. Additional details on the Gumbel-Softmax distribution and implementation of $f_{c}$ and $f_{\nu}$ are provided in Appendix B.1. To the best of our knowledge, we are the first to implement early exit using the Gumbel-Softmax reparametrization trick. Following Finkelshtein et al. (2024), for node classification, we design $f_{c}$ and $f_{\nu}$ as GNNs to estimate $\mathbf{C}^{t}$ and $\bm{\nu}^{t}$. This choice allows each node to incorporate information from its $L_{f}$-hop neighborhood—effectively reasoning $L_{f}$-steps ahead when deciding whether to exit or adjust the temperature. We observe that the decision-making process is based on the context because of the hidden features, but also depending on the task. Indeed, the update of the weights of $f_{c}$ and $f_{\nu}$ is exclusively based on the gradients of the task loss (e.g., Cross-Entropy, Mean Squared Error), since we do not use additional losses.
Early-Exit Graph Neural Networks. As shown, SAS-GNN satisfies Goal O1. By combining it with the early-exit mechanism for Goal O2, we realize our complete framework called Early-Exit Graph Neural Networks (EEGNNs), and presented in Algorithm 1. The algorithm is designed for node classification on a single graph, but can be naturally extended to inductive settings by applying it per sample. The graph-level version is provided in Appendix D. To integrate our Early-Exit mechanism into SAS-GNN, we need to include the Gumbel-Softmax scores in the message-passing update. Finkelshtein et al. (2024) implements this by modifying the graph topology, meaning that the gradients of the task loss are computed w.r.t. $\mathbf{\bar{A}}$. In our case, we decided to use the integration constant $\tau$, which is no longer fixed, but it varies across the layers and also node-wise. This approach is novel in the literature, and we refer to it as Neural Adaptive-step Early-Exit. The idea is not only to make everything differentiable but, in general, by choosing $\bm{\tau}^{l}=\mathbf{c}^{l}(\texttt{0})$ as the probability of non-exiting, we keep updating the next node representation proportionally to this probability. In fact, if $\mathbf{c}^{l}(\texttt{0})=0$, then $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}$, which enforces further the early-exit bias. At each iteration, we save those nodes that are predicted to exit. At the end of the main loop, we use the last representations $\mathbf{h}_{i}^{L}$ to complete the output. In this algorithm, we have a different integration constant per node; for this reason, we view $l$ as the $l$-th exit point. $L$ can be seen as the number of candidate exit points. In the time domain, the total time that a node $u$ spends in the GNN is computed as $\sum_{l=0}^{L}\tau^{l}_{u}$, where each $\tau^{l}_{u}$ is predicted by a GNN and also corresponds with the non-exiting probability. In this sense, we can describe the exit of each node as continuous rather than limited to the discrete domain. While this remains beyond the scope of the paper and we do not investigate this aspect extensively, in Appendix F.11, we depict how each node can exit at a unique instant of time.
To summarize, using SAS-GNN in line 6 of the algorithm offers two key benefits: (1) it provides a fallback that mitigates message-passing failures at any depth; and (2) weight sharing avoids parameter waste from unused layers.
Relation to Adaptive-Step in Runge-Kutta Solvers.Our Neural Adaptive-Step shares conceptual similarities with ODE solvers like DOPRI5/8 (Dormand and Prince, 1980), which also relax fixed-depth constraints. However, our approach differs in that a neural network determines the stopping condition, whereas DOPRI solvers rely on integration error tolerance. Although integrating these methods may be beneficial, it is outside the scope of this work.
Mitigating Under-reaching. EEGNN is inherently robust to under-reaching. We observe that setting a large $L$ does not hurt performance, as the model learns to exit early when appropriate. While related work learns depth during training (Errica et al., 2023), integrating this into our setting is left for future exploration.
Discussion on Complexity. We analyze the space complexity in terms of parameter count, comparing MPNNs, GTs, and Co-GNNs with our models. Thanks to weight sharing, SAS-GNN maintains constant complexity w.r.t. $L$ and quadratic complexity in the hidden dimension $m^{\prime}$, as both $\mathbf{\Omega}_{as}$ and $\mathbf{W}_{s}$ are $m^{\prime 2}$ matrices. Their symmetry and antisymmetry reduce the number of unique parameters to approximately $\frac{1}{2}m^{\prime 2}\cdot 2$. EEGNN adds parameters via $f_{c}$ and $f_{\nu}$, but this overhead depends only on $L_{f}$, not $L$, since these modules are shared across layers. When implemented as SAS-GNNs, the overall complexity becomes $\mathcal{O}(m^{\prime 2}+m_{f}^{2})$. Table 1 also includes Polynormer, a GT-based model that uses two modules: local attention (with $L_{l}$ layers) and global attention (with $L_{g}$ layers), each using 4 non-shared weight matrices of size $m^{\prime 2}$. Assuming $L_{l}=L_{g}=L$, the total complexity becomes $\mathcal{O}(8Lm^{\prime 2})$ as shown in table. Regarding time complexity, SAS-GNN offers no clear runtime advantage, but EEGNN may reduce computation by exiting early for many nodes. Although some nodes may still require full-depth processing, potentially becoming bottlenecks, this is not always the case, as explored in Section 4.1.
Table 1: Space complexity comparison among models.
Models GCN SAS-GNN Co-GNN EEGNN Polynormer
$\mathcal{O}(\cdot)$ $\mathcal{O}(Lm^{\prime 2})$ $\mathcal{O}(m^{\prime 2})$ $\mathcal{O}(Lm^{\prime 2}+2L_{f}m^{\prime 2}_{f})$ $\mathcal{O}(m^{\prime 2}+2L_{f}m^{2}_{f})$ $\mathcal{O}(8Lm^{\prime 2})$
4 Experimental Evaluation
In this section, we empirically validate the contributions of our proposed framework, with a particular emphasis on the multiple advantages offered by EEGNN, in terms of efficiency, effectiveness, modularity, and flexibility.
4.1 Impact of the Early-Exit Components
According to Algorithm 1, EEGNN, and its exit module (i.e., $f_{c}$, $f_{\nu}$) updates its weights directly through the task loss, conversely from previous methods, which typically require additional loss terms or regularizers to guide the learning process (Spinelli et al., 2021). This feature enables EEGNN to learn to exit in a task-driven fashion, providing a significant advantage over fixed-depth models, which otherwise require manual tuning of the optimal number of layers $L$ for each dataset and task.
To assess this, we evaluate on Peptides-func (graph classification) and Peptides-struct (graph regression), which share graph structures but differ in predictive goals (see Table 7 in the appendix), potentially requiring different problem radii. As shown in Figure 1, EEGNN effectively learns distinct exit distributions for each task. Notably, for Peptides-func, EEGNN predicts to exit after just two layers, suggesting long-range interactions are not essential for this task.
Our approach is also modular, owing to the use of the Straight-Through Gumbel-Softmax Estimator. By simply replacing line 6 in Algorithm 1, the early-exit mechanism can be integrated into other message-passing architectures. Here, parameter sharing is desirable for space efficiency, but it is not a hard constraint. To demonstrate this flexibility, in Appendix F.6, we attach the early-exit module to standard MPNNs and evaluate the combination across three datasets.
Results show that baseline models often degrade as depth increases, likely due to over-smoothing or over-squashing, while early exits help mitigate these issues, preserving or even improving performance, similarly to Spinelli et al. (2021). This highlights how depth, when not properly stabilized (e.g., through architectural choices like SAS-GNN), can hinder predictive accuracy. Later in this section, we show that equipping an already robust backbone like SAS-GNN with EEGNN does not aim to improve accuracy, but rather to enhance inference efficiency. We also find that replacing ReLU with a ReLU+TanH combination can help reduce degradation in some models.
Another substantial benefit of EEGNN is its ability to reduce inference time by skipping redundant computation for nodes that exit early. As shown in Table 2, we fix $m^{\prime}=32$ and vary $L$, averaging inference time over 1,500 forward passes. EEGNN and SAS-GNN maintain a constant number of parameters due to shared weights, unlike models such as GCN, Co-GNN, and Polynormer. While Co-GNN and Polynormer demonstrate strong performance on several benchmarks (Luo et al., 2024), they incur significantly higher computational and memory costs—Co-GNN due to expensive propagation mechanisms, and Polynormer due to its large parameter count. According to this analysis, EEGNN treats $L$ as a budget rather than a hyperparameter and effectively maintains nearly constant inference complexity. For instance, increasing $L$ does not involve a significant change to the inference time. For a more detailed runtime analysis, refer to Appendix F.9.
Figure 1: EEGNN’s Graph exit layer distributions. Numbers on top of the bars correspond to the number of test graphs that were processed up to that layer. The associated performance is in Table 4.
Table 2: Runtime and parameter analysis in Amazon Ratings. Times are in seconds.
Model Inference Time (s) Number of Parameters 10 Layers 20 Layers 10 Layers 20 Layers GCN $0.0269\pm 0.0113$ $0.0370\pm 0.0140$ 20,352 30,912 Co-GNN $0.0562\pm 0.0163$ $0.0838\pm 0.0262$ 35,049 55,849 Polynormer $0.0191\pm 0.0037$ $0.0310\pm 0.0046$ $47,306$ $80,266$ SAS-GNN $0.0278\pm 0.0106$ $0.0442\pm 0.0150$ 11,904 11,904 EEGNN $0.0267\pm 0.0128$ $0.0227\pm 0.0092$ 14,146 14,146
Table 3: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. Scores are from Finkelshtein et al. (2024); Luo et al. (2024).
Model Amazon Ratings Minesweeper Roman Empire Tolokers Questions ACC $\uparrow$ AUROC $\uparrow$ ACC $\uparrow$ AUROC $\uparrow$ AUROC $\uparrow$ GCN 53.80 ± 0.60 97.86 ± 0.52 91.27 ± 0.20 83.64 ± 0.67 79.02 ± 1.27 SAGE 55.40 ± 0.21 97.77 ± 0.62 91.06 ± 0.67 82.43 ± 0.44 77.21 ± 1.28 GAT 55.54 ± 0.51 97.73 ± 0.73 90.63 ± 0.14 83.78 ± 0.43 77.95 ± 0.51 GT 51.17 ± 0.66 91.85 ± 0.76 86.51 ± 0.73 83.23 ± 0.64 77.95 ± 0.68 Polynormer 54.81 ± 0.49 97.46 ± 0.36 92.55 ± 0.37 85.91 ± 0.74 78.92 ± 0.89 Co-GNN 54.17 ± 0.37 97.31 ± 0.41 91.37 ± 0.35 84.45 ± 1.17 76.54 ± 0.95 SAS-GNNnoedge 51.47 ± 0.68 93.29 ± 0.61 83.46 ± 0.61 85.80 ± 0.79 79.60 ± 1.15 EEGNN 51.47 ± 0.51 93.18 ± 1.37 80.36 ± 0.43 85.26 ± 0.65 78.90 ± 1.15
Table 4: LRGB datasets. The scores are marked in red for the first, blue for the second, and green for the third.
Model Peptides-func Peptides-struct Pascal VOC-SP AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$ Classic MPNNs GCN 68.60±0.50 0.2460±0.0013 20.78±0.31 GINE 66.21±0.67 0.2473±0.0017 27.18±0.54 GatedGCN 67.65±0.47 0.2477±0.0009 38.80±0.40 GTs GT+LapPE 63.26±1.26 0.2529±0.0016 26.94±0.98 SAN+LapPE 63.84±1.21 0.2683±0.0043 32.30±0.39 GraphGPS+LapPE 65.35±0.41 0.2500±0.0005 37.48±1.09 Asynchronous MPNNs Co-GNNs 69.90 ± 0.93 - - AMP 71.63 ± 0.58 0.2431 ± 0.0004 - Ours SAS-GNNnoedge 69.71 ± 0.62 0.2449 ± 0.0013 22.65 ± 0.27 SAS-GNNedge 69.27 ± 0.58 0.2547 ± 0.0163 23.31 ± 0.49 SAS-GNNours 69.44 ± 0.63 0.2528 ± 0.0130 25.64 ± 0.63 EEGNNours 68.23 ± 0.37 0.2532 ± 0.0050 24.10 ± 0.73
4.2 Comparison with the Heterophilic and LRGB Benchmarks
We evaluate EEGNN on two widely studied benchmark settings. To isolate the impact of the early-exit mechanism, we also study SAS-GNN as an ablation of EEGNN with a fixed-depth architecture. For instance, SAS-GNN required tuning the number of layers $L$ for each dataset.
Datasets. We consider two evaluation settings. First, we use the heterophilic node classification benchmarks from Platonov et al. (2023). Second, we evaluate on three LRGB datasets: Peptides-func (graph classification), Peptides-struct (graph regression), and Pascal-VOC (inductive node classification). Additional results on transductive homophilic datasets and short-range graph classification tasks from the TUDataset collection (Morris et al., 2020) are presented in Appendix F.7 and F.8. Full dataset and implementation details are available in Appendix E.
Baselines. For the heterophilic benchmarks, we follow Platonov et al. (2023), comparing against GCN, GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GT (Dwivedi and Bresson, 2021a)—simple models that have proven effective even in heterophilic settings. We also include Co-GNN, a representative asynchronous MPNN, and Polynormer, a graph Transformer. For the LRGB benchmarks, we group baselines into: (i) classical MPNNs such as GCN, GINE (Hu et al., 2020), and GatedGCN (Bresson and Laurent, 2018); (ii) graph Transformers including GT, SAN (Kreuzer et al., 2021), and GraphGPS (Rampášek et al., 2023); and (iii) asynchronous MPNNs such as Co-GNN and AMP.
Results. EEGNN and SAS-GNN, implemented as in Algorithm 1 without edge features and under Corollary 3.3, perform competitively across both heterophilic datasets (Table 3) and LRGB benchmarks (Table 4). Despite their simplicity, lacking normalization or dropout, SAS-GNN ranks in the top two on 2 out of 5 datasets, while EEGNN closely matches its performance, demonstrating that early exits incur minimal accuracy trade-off.
[…]
4.3 Empirical Validation of Theoretical Properties
Having evaluated the practical benefits of EEGNN and SAS-GNN in real-world tasks, we now turn to validating the theoretical claims underlying the SAS-GNN block, specifically, its ability to mitigate OST and OSQ according to the theoretical properties expressed in Section 3.
Over-smoothing. To assess OST, we monitor the Dirichlet Energy $E^{dir}(\mathbf{H})$, which quantifies feature smoothness over a graph:
$E^{dir}(\mathbf{H})=\sum_{(i,j)\in\mathcal{E}}\|(\nabla\mathbf{H})_{ij}\|^{2},\quad(\nabla\mathbf{H})_{ij}:=\frac{\mathbf{h}_{j}}{\sqrt{d_{j}+1}}-\frac{\mathbf{h}_{i}}{\sqrt{d_{i}+1}}.$ (5)
Lower values of $E^{dir}$ indicate smoothing; values near zero signal OST (Cai and Wang, 2020). In Figure 2(a), we report $E^{dir}(\mathbf{H}^{t})$ at each layer on Minesweeper, a heterophilic dataset. GCN rapidly collapses, while GRAFF oversharpens, as described in Di Giovanni et al. (2023). SAS-GNN and A-DGN, in contrast, maintain stable energy profiles, suggesting SAS-GNN inherits its stabilizing behavior primarily from A-DGN.
Over-squashing. We evaluate OSQ using layer-wise sensitivity $S_{l}$, defined as $S_{l}=\sum_{(v,u)\in\mathcal{E}}\left\lVert\frac{\partial\mathbf{h}_{v}^{L}}{\partial\mathbf{h}_{u}^{l}}\right\rVert_{1}$, which measures intermediate embeddings affects final representations. Due to its computational cost, we compute $S_{l}$ on one test graph. Results on Peptides-func (Figure 2(b)) show that GCN outputs are more affected by the latest layers, indicating limited early-layer influence. In contrast, GRAFF, A-DGN, and SAS-GNN maintain higher sensitivity in early layers, suggesting a more balanced propagation of information.
These trends affirm that SAS-GNN is effective as a proxy to contrast OST and OSQ. Further experimental details, Dirichlet trends, and sensitivity curves appear in Appendix F.2. Appendix F.1 also shows that SAS-GNN stability and non-dissipativeness allow retaining performance even with 100 layers and higher values of $m^{\prime}$ (see Table 8 in the appendix).
(a) $E^{dir}$ on Minesweeper.
[…]
5 Conclusions
We presented EEGNN, a framework that removes the fixed-depth constraint in GNNs through a differentiable early-exit mechanism at both node and graph levels. At its core lies SAS-GNN, a stable, non-dissipative message-passing scheme designed to mitigate over-smoothing and over-squashing. Empirical results across heterophilic and long-range benchmarks show that EEGNN delivers competitive performance while remaining time and memory-efficient.
Limitations. Despite these advantages, our models may underperform on tasks that demand higher expressive power, likely due to their architectural simplicity. Moreover, we do not directly address the expressiveness limitations of the 1-WL test, nor provide a principled solution to under-reaching.
Future Work. We plan to explore more expressive architectures following the example of this manuscript, by improving the early-exit strategies, and extending EEGNN to edge-level tasks such as link prediction, where exit decisions are more complex due to the pairwise nature of edge representations.
Appendix A Appendix Overview
  • Appendix B: Additional Preliminaries.
    We provide information related to GRAFF and A-DGN, which are the foundation of SAS-GNN, but are not discussed in the main paper. We also provide some preliminaries on the Gumbel-softmax distribution.
  • Appendix C: Proofs of Theorems in Section 4.
    Here, we provide the proofs of the theoretical results that support the design of SAS-GNN. We also show that including edge features into the SAS-GNN update rule, SAS-GNN preserves its stability, non-dissipativeness, and its capability of inducing attraction and repulsion.
  • Appendix D: Neural Adaptive-Step Early-Exit for Graph Classification.
    We report the extension of Algorithm 1 for node classification to the task of classifying graphs.
  • Appendix E: Additional Details on the Experimental Setup.
    This section presents additional information on the datasets, training procedures, and the hardware used in the experiments.
  • Appendix F: Additional Results.
    This section presents extended results from our experiments with EEGNN. We bring additional evidence that SAS-GNN’s design is a proxy to mitigate over-smoothing and over-squashing, and also to retain performance as depth increases. We provide additional baselines on the heterophilic benchmark and LRGB datasets. We provide results on homophilic node classification and graph classification datasets. We test how classic GNNs perform when equipped with our early-exit module. We provide evidence that Early-Exit is a promising direction to significantly improve the GNN performance. We illustrate the nodes’ exit distributions in the discrete and continuous cases that we obtain for the other datasets.
  • Appendix G: Extended Related Work.
    In the main paper, we focused our discussion on related work, mainly on GNN papers where the targets are over-smoothing and over-squashing, or papers discussing the adaptive depth choice. Here we extend the discussion by also including GNNs and Early-Exit Neural Networks in general.
Appendix B Additional Preliminaries
Anti-Symmetric Deep Graph Networks. Let a message-passing update for a node $i$ be
$\dot{\mathbf{h}}_{i}^{t}=f(\mathbf{h}_{i}^{t})=\sigma(\mathbf{\Omega}\mathbf{h}_{i}^{t}+\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))+b),$ (6)
where $\mathbf{\Omega},\mathbf{W}\in\mathbb{R}^{d\times d}$ are trainable matrices, $\sigma$ is a non-linearity, $\mathbf{h}_{i}^{t}$ is the feature of node $i$ at time $t$, $\Gamma(i)$ is the set of neighbors of $i$, $\mathbf{H}^{t}\in\mathbb{R}^{n\times d}$ is the matrix containing all the $d$ dimensional features for each node, $b\in\mathbb{R}^{d}$. The following definitions are from Gravina et al. [2023].
Definition B.1.
A solution $\mathbf{h}_{i}^{t}$ of the ODE in Equation 6, with initial condition $\mathbf{h}_{i}^{0}$, is stable if for any $\omega>0$, there exists a $\delta>0$ such that any other solution $\mathbf{\tilde{h}}_{i}^{t}$ of the ODE with initial condition $\mathbf{\tilde{h}}_{i}^{0}$ satisfying $|\mathbf{h}_{i}^{0}-\mathbf{\tilde{h}}_{i}^{0}|\leq\delta$ also satisfies $|\mathbf{h}_{i}^{t}-\mathbf{\tilde{h}}_{i}^{t}|\leq\omega$, for all $t\geq 0$.
Definition B.2.
Let $E\subseteq\mathbb{R}^{d}$ be a bounded set that contains any initial condition $\mathbf{h}_{i}^{0}$ for the ODE in Equation 6. The system defined by the ODE in Equation 6 is dissipative if there is a bounded set $B$ where, for any $E$, there exists $t^{*}\geq 0$ such that $\{\mathbf{h}_{i}^{t}\mid\mathbf{h}_{i}^{0}\in E\}\subseteq B$ for $t>t^{*}$.
In Gravina et al. [2023], the authors propose an instance of Equation (6) that is both stable and non-dissipative, leading to an evolution of node features that retain all the information collected during the forward pass (e.g., from 00 to an arbitrary $t$). Such an instantiation is
$\dot{\mathbf{h}}_{i}^{t}=f(\mathbf{h}_{i}^{t})=\text{tanh}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t}+\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))+b),$ (7)
which can be discretized through the Euler method as
$\frac{\mathbf{h}_{i}^{t+\tau}-\mathbf{h}_{i}^{t}}{\tau}=f(\mathbf{h}_{i}^{t})=\text{tanh}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t}+\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))+b).$ (8)
Here, the author uses the $\tanh$ to keep bounded the Jacobian of the system, and $\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))$, do not have any dependency on $\mathbf{h}_{i}^{t}$.
Graph Neural Networks as Gradient Flows. Another message-passing rule [Di Giovanni et al., 2023] exploits symmetric matrices $\mathbf{\Omega}_{s}$ and $\mathbf{W}_{s}$ to contrast over-smoothing in heterophilic graphs, which is a desired feature for message-passing neural networks. Using the symmetric bias, they are minimizing an energy functional that induces both attraction and repulsion to connected nodes via the eigenvalues of the weight matrices. This is the message-passing update rule:
$\mathbf{H}^{t+\tau}=\mathbf{H}^{t}+\tau\sigma(\mathbf{H}^{t}\mathbf{\Omega}_{s}+\mathbf{A}\mathbf{H}^{t}\mathbf{W}_{s}).$ (9)
As long as $\sigma$ is a non-linearity s.t. $x\sigma(x)\geq 0$ (e.g., ReLU()($\cdot$)( ⋅ ) or tanh()($\cdot$)( ⋅ )), and the weight matrices are symmetric, this rule is proved in Di Giovanni et al. [2023] to minimize the underlying functional:
$E_{\theta}^{dir}(\mathbf{H})$ $=\sum_{i}\langle\mathbf{h}_{i},\mathbf{\Omega}_{s}\mathbf{h}_{i}\rangle-\sum_{i,j}a_{ij}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{j}\rangle$ (10)
$=\sum_{i}\langle\mathbf{h}_{i},(\mathbf{\Omega}_{s}-\mathbf{W}_{s})\mathbf{h}_{i}\rangle+\sum_{i}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{i}\rangle-\sum_{i,j}a_{ij}\langle\Theta_{+}\mathbf{h}_{i},\Theta_{+}\mathbf{h}_{j}\rangle+\sum_{i,j}a_{ij}\langle\Theta_{-}\mathbf{h}_{i},\Theta_{-}\mathbf{h}_{j}\rangle$ (11)
$=\sum_{i}\langle\mathbf{h}_{i},(\mathbf{\Omega}_{s}-\mathbf{W}_{s})\mathbf{h}_{i}\rangle+\frac{1}{2}\sum_{i,j}\|\Theta_{+}(\nabla\mathbf{H})_{ij}\|^{2}-\frac{1}{2}\sum_{i,j}\|\Theta_{-}(\nabla\mathbf{H})_{ij}\|^{2},$ (12)
we consider $a_{i,j}=\frac{1}{\sqrt{d_{i}\cdot d_{j}}}$, which assigns to each edge a weight that depends on the degrees $d_{i}$, $d_{j}$ of the nodes $i$ and $j$. To arrive at this form, we leverage the symmetry of $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$, which allows spectral decomposition as $\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu})\Psi^{\top}$. The eigenvalue vector $\boldsymbol{\mu}$ can be split into its positive and negative components, yielding the decomposition:
$\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu}_{+})\Psi^{\top}+\Psi\mathrm{diag}(\boldsymbol{\mu}_{-})\Psi^{\top}=\mathbf{W}_{+}-\mathbf{W}_{-},$
where $\mathbf{W}_{+}$ and $\mathbf{W}_{-}$ are real, symmetric, and positive semi-definite matrices.
We then apply the Cholesky decomposition to each term, expressing:
$\mathbf{W}_{+}=\Theta_{+}^{\top}\Theta_{+},\qquad\mathbf{W}_{-}=\Theta_{-}^{\top}\Theta_{-},$
where $\Theta_{+},\Theta_{-}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are lower triangular matrices. Substituting these into the energy functional results in Equation (10), which clearly separates the smoothing (attraction) and sharpening (repulsion) effects. The positive semi-definite part $\Theta_{+}$ contributes to smoothing by encouraging alignment between neighboring node features, while $\Theta_{-}$ induces repulsion, preserving sharp differences. We can see that a gradient flow for such energy, namely Equation (9), can minimize or maximize the edge gradients computed on the node features. This results in a model’s behavior that allows for attraction and repulsion.
B.1 The Gumbel Distribution and the Gumbel-Softmax Temperature
The following exposition largely follows Finkelshtein et al. [2024]. The Gumbel distribution is widely used to model the maximum (or minimum) of a set of random variables. Its probability density function is asymmetric and has heavy tails, making it suitable for representing rare or extreme events. When applied to logits or scores corresponding to discrete choices, the Gumbel-Softmax estimator transforms these into a probability distribution over the available options.
The probability density function of a variable $X\sim\text{Gumbel}(0,1)$ is defined as:
$f(x)=e^{-x-e^{-x}},$ (13)
and is shown in Figure 3.
The Straight-Through Gumbel-Softmax estimator typically benefits from learning an inverse temperature parameter before sampling actions, a strategy we adopt in our experiments. For a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, the inverse temperature for node $v\in\mathcal{V}$ is computed by applying a bias-free linear layer $f_{\nu}:\mathbb{R}^{m^{\prime}}\rightarrow\mathbb{R}$ to the intermediate node representation $\mathbf{H}^{l}\in\mathbb{R}^{n\times m^{\prime}}$. To ensure that the temperature remains positive, we apply a smooth approximation of the ReLU function followed by a bias term $\nu_{0}\in\mathbb{R}$:
$f_{\nu}(\mathbf{H}^{l},\nu_{0})=\frac{1}{\bm{\nu}^{l}}=\log\left(1+\exp\left(\text{g}(\mathbf{H}^{l})\right)\right)+\nu_{0}$ (14)
Here, $\nu_{0}$ sets the minimum inverse temperature, thereby controlling the upper bound of the temperature. In our setup, we implement $f_{\nu}$ as in Equation 14, with $\text{g}=\text{GNN}(\cdot)$ for node classification (see Algorithm 1) and $\text{g}=\text{MLP}(\cdot)$ for graph classification (see Algorithm 2). The confidence estimator $f_{c}$ is defined as $\text{g}(\mathbf{H}^{l})$, choosing g based on the task, similarly to $f_{\nu}$.
Figure 3: The pdf $f(x)=e^{-x-e^{-x}}$ of Gumbel(0, 1).
Appendix C Proofs of Section 3
We want to prove and understand the conditions under which our message-passing rule is stable, non-dissipative, and can induce attraction or repulsion among adjacent nodes. Our proposal is the following:
$\dot{\mathbf{H}}^{t}=\sigma_{1}(-\sigma_{2}(\mathbf{H}^{t}\mathbf{\Omega}_{as})+\mathbf{\bar{A}}\mathbf{H}^{t}\mathbf{W}_{s}),$ (15)
where $\mathbf{\Omega}_{as}=\mathbf{\Omega}-\mathbf{\Omega}^{\top}$. We state two different Theorems to understand if we can guarantee all of these specifics. In the first proof, we leverage the node-wise Equation for $\mathbf{h}_{i}^{t}$ to demonstrate that each node evolution is stable and non-dissipative.
The following exposition share similarities with Chang et al. [2019], Gravina et al. [2023], Di Giovanni et al. [2023].
C.1 Proof of Theorem 3.1: stability and non-dissipativeness
Proof of the Theorem. We assume that we have a message-passing of the type:
$\dot{\mathbf{h}}_{i}^{t}=\sigma_{1}(-\sigma_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top}))\mathbf{h}_{i}^{t}+\phi_{\mathbf{W_{s}}}(\mathbf{H}^{t},\Gamma(i))),$ (16)
where the derivative of $\sigma_{1}$ is a bounded and point-wise non-linear function, $\sigma_{2}$ is a non-linear activation function, $\phi_{\mathbf{W_{s}}}(\mathbf{H}^{t},\Gamma(i))$ do not depend on $\mathbf{h}_{i}^{t}$, and $\mathbf{W}_{s}$ is a symmetric matrix. We can rewrite (6) as:
$\frac{\partial\mathbf{h}_{i}^{t}}{dt}=f(\mathbf{h}_{i}^{t})$ (17)
If we differentiate (17) w.r.t. the initial condition of the node $i$ we have:
$\frac{d}{d\mathbf{h}_{i}^{0}}\frac{d\mathbf{h}_{i}^{t}}{dt}=\frac{d}{\partial\mathbf{h}_{i}^{0}}f(\mathbf{h}_{i}^{t}),$ (18)
which can be adjusted through the chain rule as
$\frac{d}{dt}\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}}=\frac{d(f(\mathbf{h}_{i}^{t}))}{\partial\mathbf{h}_{i}^{t}}\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}},$ (19)
where $\frac{d(f(\mathbf{h}_{i}^{t}))}{\partial\mathbf{h}_{i}^{t}}$ is the Jacobian of the system. So we rewrite Equation (19) in these terms.
$\frac{d}{dt}\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}}=J^{t}\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}},$ (20)
Here, assuming that $J^{t}$ does not change with $t$, and $\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}}$ as the system’s state variables, we can write the solution of Equation (20) as
$\frac{\partial\mathbf{h}_{i}^{t}}{\partial\mathbf{h}_{i}^{0}}=e^{tJ}=\mathbf{T}e^{t\mathbf{\Lambda}}\mathbf{T}^{-1}$
where we apply the spectral decomposition of $J$, and get its eigenvectors $\mathbf{T}$, and eigenvalues $\mathbf{\Lambda}=diag(\lambda_{i})$. As denoted in Gravina et al. [2023], to guarantee the stability of the system, we require the $Re(\lambda_{i})\leq 0,\forall i=1,\dots,d$. However, when the eigenvalues are less than zero, we can lose the information of the previous node representations, leading to a dissipative behavior. For this reason we want $Re(\lambda_{i})=0,\forall i=1,\dots,d$, having only imaginary eigenvalues. In the case of Equation (7), the Jacobian has only imaginary eigenvalues and takes the form of
$J^{t}=\text{diag}[\text{tanh}^{\prime}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t}+\phi_{W}(\mathbf{H}^{t},\Gamma(i))+b)](\mathbf{\Omega}-\mathbf{\Omega}^{\top})$ (21)
Here, the choice of the antisymmetric multiplication by $\mathbf{\Omega}-\mathbf{\Omega}^{\top}$ guarantees the presence of imaginary eigenvalues, and left-multiplying it by a diagonal matrix does not change the nature of the overall spectrum. This was proved in Gravina et al. [2023]. Our Jacobian instead becomes:
$J^{t}=-\text{diag}[\sigma_{1}^{\prime}(-\sigma_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})+\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i)))]\text{diag}[\sigma^{\prime}_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})](\mathbf{\Omega}-\mathbf{\Omega}^{\top}).$ (22)
Since we assume that the derivative of $\sigma_{1}$ is bounded, the assumption that $J^{t}$ stays constant can be valid. Hereafter, we show this proof, similarly to Gravina et al. [2023].
Defining $\mathbf{A}=-\text{diag}[\sigma_{1}^{\prime}(-\sigma_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})+\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i)))]\text{diag}[\sigma^{\prime}_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})]$ and $\mathbf{B}=(\mathbf{\Omega}-\mathbf{\Omega}^{\top})$, we have:
$J^{t}=\mathbf{AB}.$
We now want to prove that the eigenvalues of $\mathbf{AB}$ are purely imaginary, namely $Re(\lambda_{i})=0$ for all $\lambda_{i}$ eigenvectors of $\mathbf{AB}$. If all eigenvectors are zero, then the statement holds. Let us now consider an eigenpair of AB, where the eigenvector is denoted by $\mathbf{v}$ and the eigenvalue by $\lambda$. Then:
$\mathbf{ABv}=\lambda\mathbf{v},$ (23)
$\mathbf{A}$ is a diagonal matrix with non-negative entries, let us denote by $\mathcal{I}\in[d]$ the set of indices s.t. =$\mathcal{I}=\{i\in[d]:\mathbf{A}_{ii}=0\}$. If $\mathbf{v}$ eigenvector of $\mathbf{AB}$ then $\mathbf{v}_{j}=0\;\forall j\in\mathcal{I}$. Indeed, we need for each entry $k$, $\sum_{i}\mathbf{AB}_{ki}v_{i}=\lambda v_{k}$; for $j$ in $\mathcal{I}$ the row the $j-$th row of $\mathbf{AB}$ is 0. So we have $\lambda v_{j}=0$, since we assumed $\lambda\neq 0,$ this implies $v_{j}=0$.
Let $\mathbf{D}$ be diagonal matrix with entries that are defined as $\mathbf{D}_{ii}=\sqrt{\mathbf{A}_{ii}}$. We moreover denote by $\tilde{\mathbf{D}}^{-1}$ the matrix defined as
$\tilde{\mathbf{D}}^{-1}_{ii}=\begin{cases}\frac{1}{\mathbf{D}_{ii}}\;\;\text{if}\;i\notin\mathcal{I}\\ 0\;\;\text{if}\;i\in\mathcal{I}.\end{cases}$ (24)
and by $\tilde{\mathbf{I}}$ the diagonal matrix defined as
$\tilde{\mathbf{I}}_{ii}=\begin{cases}1\;\;\text{if}\;i\notin\mathcal{I}\\ 0\;\;\text{if}\;i\in\mathcal{I}.\end{cases}$ (25)
Notice that $\tilde{\mathbf{D}}^{-1}{\mathbf{D}}={\mathbf{D}}\tilde{\mathbf{D}}^{-1}=\tilde{\mathbf{I}}$, and that $\tilde{\mathbf{I}}\mathbf{A}=\mathbf{A}$ and $\tilde{\mathbf{I}}\mathbf{D}=\mathbf{D}$.
Let $\mathbf{M}$ be defined as $\mathbf{M}=\tilde{\mathbf{D}}^{-1}\mathbf{AB}\mathbf{D}$.
$\mathbf{M}$ is skew symmetric, indeed $\mathbf{M}=\tilde{\mathbf{D}}^{-1}\mathbf{AB}\mathbf{D}=\tilde{\mathbf{D}}^{-1}\mathbf{DDB}\mathbf{D}=\tilde{\mathbf{I}}\mathbf{DBD}=\mathbf{DBD}$.
It is straightforward to derive that $\mathbf{M}$ is so that the rows in the set $\mathcal{I}$ are null vectors.
We have that for every vector $\mathbf{w}$ s.t. $\mathbf{w}_{j}=0\;\forall j\in\mathcal{I}$, it holds that
$\mathbf{DM}\tilde{\mathbf{D}}^{-1}\mathbf{w}=\mathbf{ABw}.$ (26)
This property can be easily verified by noticing that if $\mathbf{w}$ has such a property, then $\tilde{\mathbf{I}}\mathbf{w}=\mathbf{w}$. Doing all the calculations, $\mathbf{DM}\tilde{\mathbf{D}}^{-1}\mathbf{w}=\mathbf{D\tilde{\mathbf{D}}^{-1}}\mathbf{ABD}\tilde{\mathbf{D}}^{-1}\mathbf{w}=\tilde{\mathbf{I}}\mathbf{AB}\tilde{\mathbf{I}}\mathbf{w}=\mathbf{AB}\mathbf{w}.$
We can now prove that every eigenvector of $\mathbf{AB}$ is also an eigenvector of $\mathbf{M}$.
We have already observed that if $\mathbf{v}$ eigenvector of $\mathbf{AB}$, it has entries in $\mathcal{I}$ equal to 00. If $\mathbf{AB}\mathbf{v}=\lambda\mathbf{v}$, using (26),
$\mathbf{DM}\tilde{\mathbf{D}}^{-1}\mathbf{v}=\lambda\mathbf{v}$, from which using that all the rows in $\mathbf{M}$ with index in $\mathcal{I}$ are null vectors and that similarly the entries of $\mathbf{v}$ with index in $\mathcal{I}$ are null we obtain, $\mathbf{M}\tilde{\mathbf{D}}^{-1}\mathbf{v}=\lambda\tilde{\mathbf{D}}^{-1}\mathbf{v}$.
Since $\mathbf{M}$ is skew-symmetric, it has only pure imaginary eigenvalues, so $\lambda$ is pure imaginary.
As a result, all eigenvalues of $\mathbf{J}^{t}$ are purely imaginary.
As we can see, this process remains the same regardless of the symmetric matrix $\mathbf{W}_{s}$. We must assure that $\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))$ does not depend on $\mathbf{h}_{i}$, otherwise the Jacobian would have become
$J^{t}=-\text{diag}[\sigma_{1}^{\prime}(-\sigma_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})+\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i)))]\text{diag}[\sigma^{\prime}_{2}((\mathbf{\Omega}-\mathbf{\Omega}^{\top})\mathbf{h}_{i}^{t})](\mathbf{\Omega}-\mathbf{\Omega}^{\top})+C),$ (27)
we can guarantee that there is no dependency, since we consider $\phi_{\mathbf{W}}(\mathbf{H}^{t},\Gamma(i))=\bar{\mathbf{A}}\mathbf{H}^{t}\mathbf{W}_{s}$, with no self-loops in $\bar{\mathbf{A}}$. Thus, we can conclude that Equation 16 is stable and non-dissipative, even if $\mathbf{W}_{s}$ is symmetric, and no specific characteristics are associated with $\sigma_{2}$; it must only be a point-wise function. Moreover, the derivative of $\sigma_{1}$ must be bounded within an interval, in order to let the constant Jacobian assumption be valid. This concludes the proof of the Theorem.
C.2 Proof of the Theorem 3.2: attraction and repulsion
Now, what we aim to understand is whether Equation 15 can also induce attraction and repulsion edge-wise as in Di Giovanni et al. [2023]. Our assumptions are the same as the previous proof, but we also assume $\sigma_{2}$ to be a point-wise activation function that returns only positive values, which do not violate the assumption of the previous proof.
Taking into consideration Equation (10), we notice that $\mathbf{\Omega}$ is not the main responsible for the attraction and repulsion behavior, but is $\mathbf{W}_{s}$. Because of this, we want to understand whether substituting $\mathbf{\Omega_{s}}$ with $\mathbf{\Omega_{as}}$ could preserve stability, non-dissipativeness, edge-wise attraction and repulsion. By simply doing this substitution, we have:
$E_{\theta}^{dir}(\mathbf{H})$ $=\sum_{i}\langle\mathbf{h}_{i},\mathbf{\Omega}_{as}\mathbf{h}_{i}\rangle-\sum_{i,j}a_{ij}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{j}\rangle$ (28)
$=\sum_{i}\langle\mathbf{h}_{i},(\mathbf{\Omega}_{as}-\mathbf{W}_{s})\mathbf{h}_{i}\rangle+\sum_{i}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{i}\rangle-\sum_{i,j}a_{ij}\langle\Theta_{+}\mathbf{h}_{i},\Theta_{+}\mathbf{h}_{j}\rangle+\sum_{i,j}a_{ij}\langle\Theta_{-}\mathbf{h}_{i},\Theta_{-}\mathbf{h}_{j}\rangle$ (29)
$=\sum_{i}\langle\mathbf{h}_{i},(\mathbf{\Omega}_{as}-\mathbf{W}_{s})\mathbf{h}_{i}\rangle+\frac{1}{2}\sum_{i,j}\|\Theta_{+}(\nabla\mathbf{H})_{ij}\|^{2}-\frac{1}{2}\sum_{i,j}\|\Theta_{-}(\nabla\mathbf{H})_{ij}\|^{2}.$ (30)
which seems to imply that the antisymmetry would not affect the attraction/repulsion framework. However, the gradient flow of this energy would become:
$\dot{\mathbf{H}}^{t}=(\mathbf{H}^{t}\mathbf{\Omega}_{as}+\mathbf{H}^{t}\mathbf{\Omega}_{as}^{\top})+\mathbf{A}\mathbf{H}^{t}\mathbf{W}_{s}$ (31)
which leads to:
$\dot{\mathbf{H}}^{t}=\mathbf{A}\mathbf{H}^{t}\mathbf{W}_{s},$ (32)
due to the antisymmetric matrix properties of $\mathbf{\Omega}_{as}=-\mathbf{\Omega}_{as}^{\top}$. Unfortunately, Equation 32 would not exploit the antisymmetric properties associated with the Jacobian, and the results of the previous proof would not be valid.
So, we cannot directly use antisymmetric weights to minimize the energy in Equation (10), since antisymmetric matrices do not yield a positive semi-definite quadratic form. However, we can still investigate whether the message-passing dynamics defined in Equation (15) induce edge-wise attraction and repulsion. To do so, we define an alternative energy functional whose minimization would have equivalent edge-level effects.
We consider the following energy:
$E_{\theta}(\mathbf{H})$ $=-\sum_{i,j}a_{ij}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{j}\rangle$ (33)
$=-\sum_{i}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{i}\rangle+\sum_{i}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{i}\rangle-\sum_{i,j}a_{ij}\langle\Theta_{+}\mathbf{h}_{i},\Theta_{+}\mathbf{h}_{j}\rangle+\sum_{i,j}a_{ij}\langle\Theta_{-}\mathbf{h}_{i},\Theta_{-}\mathbf{h}_{j}\rangle$ (34)
$=-\sum_{i}\langle\mathbf{h}_{i},\mathbf{W}_{s}\mathbf{h}_{i}\rangle+\frac{1}{2}\sum_{i,j}\|\Theta_{+}(\nabla\mathbf{H})_{ij}\|^{2}-\frac{1}{2}\sum_{i,j}\|\Theta_{-}(\nabla\mathbf{H})_{ij}\|^{2},$ (35)
where $(\nabla\mathbf{H})_{ij}=\frac{\mathbf{h}_{i}}{\sqrt{d_{i}}}-\frac{\mathbf{h}_{j}}{\sqrt{d_{j}}}$ is the discrete node-wise gradient.
To arrive at this form, we leverage the symmetry of $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$, which allows spectral decomposition as $\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu})\Psi^{\top}$. The eigenvalue vector $\boldsymbol{\mu}$ can be split into its positive and negative components, yielding the decomposition:
$\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu}_{+})\Psi^{\top}+\Psi\mathrm{diag}(\boldsymbol{\mu}_{-})\Psi^{\top}=\mathbf{W}_{+}-\mathbf{W}_{-},$
where $\mathbf{W}_{+}$ and $\mathbf{W}_{-}$ are real, symmetric, and positive semi-definite matrices.
We then apply the Cholesky decomposition to each term, expressing:
$\mathbf{W}_{+}=\Theta_{+}^{\top}\Theta_{+},\qquad\mathbf{W}_{-}=\Theta_{-}^{\top}\Theta_{-},$
where $\Theta_{+},\Theta_{-}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are lower triangular matrices. Substituting these into the energy functional results in Equation (35), which clearly separates the smoothing (attraction) and sharpening (repulsion) effects. The positive semi-definite part $\Theta_{+}$ contributes to smoothing by encouraging alignment between neighboring node features, while $\Theta_{-}$ induces repulsion, preserving sharp differences.
Although this energy is no longer a direct generalization of the classic Dirichlet energy, the spectral properties of $\mathbf{W}_{s}$ still allow edge-level attraction and repulsion to emerge via its eigenstructure. In the following, we analyze the time derivative of this energy using the dynamics from Equation (15), to verify whether the system implicitly minimizes it through evolution.
$\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t})=-\mathbf{A}\mathbf{H}^{t}(\frac{\mathbf{W}_{s}+\mathbf{W}_{s}}{2})=-\mathbf{A}\mathbf{H}^{t}\mathbf{W}_{s},$ (36)
Without loss of generality, we use a unique symmetric matrix $\mathbf{W}_{s}$. If we want to represent $\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t})$ in a vectorized form, where we stack all the columns together, $\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t})\in\mathbb{R}^{n\times m^{\prime}}$ becomes $\text{vec}(\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t}))\in\mathbb{R}^{nm^{\prime}\times 1}$. According to this we derive Equation (36) through the kronecker product $\otimes$ formalism as
$\text{vec}(\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t}))=-(\mathbf{W}_{s}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{H}^{t})$ (37)
And we can also derive Equation (15) as
$\text{vec}(\dot{\mathbf{H}}^{t})=\sigma_{1}((\mathbf{W}_{s}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{H}^{t})-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t})))$ (38)
The time derivative through this formalism can be written as:
$\frac{dE_{\theta}(\mathbf{H}^{t})}{dt}$ $=\text{vec}(\nabla_{\mathbf{H}}E_{\theta}(\mathbf{H}^{t}))^{\top}\text{vec}(\dot{\mathbf{H}}^{t})$ (39)
$=-(\mathbf{W}_{s}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{H}^{t})\sigma_{1}((\mathbf{W}_{s}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{H}^{t})-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t})))$ (40)
$=-\mathbf{Z}^{t}\sigma(\mathbf{Z}^{t}+\mathbf{Y}^{t})$ (41)
We consider $\mathbf{Z}^{t}=(\mathbf{W}_{s}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{H}^{t})$ and $\mathbf{Y}^{t}=-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t}))$, which are $nm^{\prime}$-dimensional vectors. We refer to $\mathbf{Z}^{t}_{i}$ and $\mathbf{Y}^{t}_{i}$ as their $i$-th component. Every choice of $\sigma$ s.t. the following condition is verified
$\mathbf{Z}^{t}_{i}\cdot\sigma(\mathbf{Z}^{t}_{i}+\mathbf{Y}^{t}_{i})\geq 0\hskip 14.22636pt\forall i=1,...nm^{\prime},$ (42)
will make Equation (39) monotonically decreasing as $t\rightarrow\infty$, and the associated evolution equation would be expressive enough to induce attraction and repulsion edge-wise. What would be such a choice of non-linearity?
ReLU, or $tanh$, does not satisfy this condition, thus, we elaborate on this step by imposing the following constraints.
$\frac{dE_{\theta}(\mathbf{H}^{t})}{dt}=-\mathbf{Z}^{t}\sigma(\mathbf{Z}^{t}-\text{ReLU}(\mathbf{Y}^{t}))$ (43)
$\frac{dE_{\theta}(\mathbf{H}^{t})}{dt}=-\mathbf{Z}^{t}\text{ReLU}(\text{tanh}((\mathbf{Z}^{t}-\text{ReLU}(\mathbf{Y}^{t}))$ (44)
The conditions in Equation (42) are always satisfied, since when $\mathbf{Z}^{t}_{i}<0$, we have that everything is 00. Otherwise, we consider two cases, namely when $\mathbf{Z}^{t}_{i}>\text{ReLU}(\mathbf{Y}^{t}_{i})$ and $\mathbf{Z}^{t}_{i}<\text{ReLU}(\mathbf{Y}^{t}_{i})$. In the former, we have that $\text{ReLU}(\text{tanh}(\cdot))$ is always positive, and also $\mathbf{Z}^{t}_{i}$. In the latter case, we have a negative argument, and the Equation goes to 00. Generally, we could satisfy these conditions for any choice of $\sigma_{1}(x)\geq 0,\forall x\in\mathbb{R}$, and also for $\sigma_{2}\geq 0,\forall x\in\mathbb{R}$. We conclude that through our assumptions, Equation (15) can induce attraction and repulsion edge-wise.
C.3 Proof of Corollary 3.3
Assuming that $\sigma_{1}=\text{ReLU}(\text{tanh}(\cdot))$, $\sigma_{2}=\text{ReLU}(\cdot)$, and considering $\bar{\mathbf{A}}$, with no self-loops, we can enclose all the assumptions made in our Theorems’ proofs. Such an instantiation of $\sigma_{1}$ would have a bounded derivative that guarantees the Jacobian being approximately constant as follows:
$\sigma^{\prime}_{1}(x)=(\text{ReLU}(\text{tanh}(x)))^{\prime}\begin{cases}1-\tanh^{2}(x)&\text{if }x>0\\ 0&\text{if }x\leq 0\end{cases},$ (45)
then the time derivative of the energy functional will be monotonically decreasing, and the Jacobian of each node feature will be constant and with purely imaginary eigenvalues. These steps are easy to verify once the proofs of the Theorems have been understood. According to these assumptions, we conclude that Equation (15), with $\sigma_{1}=\text{ReLU}(\text{tanh}(\cdot))$, $\sigma_{2}=\text{ReLU}(\cdot)$, is stable and non-dissipative, and induces attraction and repulsion edge-wise.
C.4 Proof of Theorem 3.4
Proof of the theorem. Theorem 3.4, is simple to demonstrate, since $-\sigma_{2}(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$ is not dependent on the node features and can be encompassed in the $\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i))$ term of Equation (16), and repeat the same proof. According to this, Equation (4) is stable and non-dissipative. To prove that it induces attraction and repulsion edge-wise, we can take Equation (39), and now specify that $\mathbf{Y}^{t}=-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t}))-\sigma_{2}((\mathbf{E}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{B}))$. We will derive the same result, since each element of $\mathbf{Y}^{t}$ is less than or equal to 0.
This concludes the proof.
Appendix D Early-Exit Graph Neural Networks Algorithm for Graph Classification
In Algorithm 1, we describe how the Straight-Through Gumbel-Softmax estimator is integrated with SAS-GNN to allow each node to decide, at every layer, whether to exit or continue processing. However, since we also address graph classification, we refine this procedure to enable early-exit decisions at the graph level. Specifically, instead of applying the Gumbel-Softmax to individual nodes as in Algorithm 1, we apply it to the entire graph—so that the model determines whether a graph’s representation is sufficiently complete to make a prediction. This modification ensures that exit decisions align with the prediction target: nodes for node classification, graphs for graph classification. To maintain consistency, we define the "agent"—the entity responsible for triggering the exit—as the unit on which the prediction task is performed: nodes for node-level tasks and graphs for graph-level tasks. While it is theoretically possible to perform graph classification with node-level exits (e.g., by aggregating node predictions), we leave such extensions for future work. The refined algorithm follows.
Algorithm 2 Neural Adaptive-Step Early-Exit GNNs for Graph Classification
1:Initialize $\mathcal{G}=(\mathbf{H}^{0},\mathbf{\bar{A}}),\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d},L,f_{c},f_{\nu},f_{e},\text{Pool},\mathbf{W}_{s},\mathbf{\Omega}_{as},\sigma_{1},\sigma_{2},\nu_{0}$
2:for $l=0$ to $L$ do
3:     $C^{l}\leftarrow f_{c}(\text{Pool}(\mathbf{H}^{l}))$
4:     $\nu^{l}\leftarrow f_{\nu}(\text{Pool}(\mathbf{H}^{l},\nu_{0}))$
5:     $\mathbf{c}^{l}\leftarrow gumbel\_softmax(C^{l},\nu^{l})$
6:     $\tau^{l}\leftarrow\mathbf{c}^{l}(\texttt{0})$
7:     $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}+\tau^{l}\sigma_{1}(-\sigma_{2}(\mathbf{H}^{l}\mathbf{\Omega}_{as})+f_{e}(\mathbf{E})+\mathbf{\bar{A}}\mathbf{H}^{l}\mathbf{W}_{s})$
8:     if $\arg\max\{\mathbf{c}^{l}\}=\texttt{1}$ then
9:         $\mathbf{Z}\leftarrow pool(\mathbf{H}^{l})$
10:         return $\mathbf{Z}$
11:$\mathbf{Z}\leftarrow pool(\mathbf{H}^{L})$
12:return $\mathbf{Z}$
Here, we consider the case of a single graph, but it can be iterated all over the dataset $\mathcal{D}$ as well. As we can see, $C^{l}$ and $\nu^{l}$ are scalars, conversely from node classification, since we have pooled the node representation. As a consequence, $\tau^{l}$ is a scalar, and is not defined node-wise anymore. Despite this, we retain the same meaning, indeed $\tau^{l}=0$ leads the algorithm to stop updating the node features ($\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}$). We understand that now the personalization will be at the graph level. Indeed for graphs $i$ and $j$, we will have $\tau_{i}^{l}\neq\tau_{j}^{l}$, since these depends on $\mathbf{H}^{l}_{i}$ and $\mathbf{H}^{l}_{j}$ respectively. In the same fashion as Algorithm 1, we have a varying integration constant, namely $\tau^{l}$. This value is continuous, but we denote $l$ as the $l$-th exit point. Also, here $L$ can be seen as the number of candidate exit points. In the time domain, the total time that a graph undergoes into the GNN is computed as $\sum_{l=0}^{L}\tau^{l}$, where each $\tau^{l}$ is predicted by $f_{c}$ and $f_{\nu}$ and it corresponds with the non-exiting probability.
Appendix E Additional Details on the Experimental Set-Up
E.1 Datasets
Table 5 lists some statistics of the heterophilic benchmark that we use. They are computed after the graphs were made undirected, which is a standard procedure with GNNs. The datasets selected belong to a recent collection [Platonov et al., 2023], which was proposed to enrich the current dataset availability for the GNN experimental setting under heterophily. $\xi_{edge}$ (edge homophily) and $\xi_{adj}$ (adjusted homophily) are metrics that estimate the heterophily level of a graph; the lower they are, the more the graph is considered as heterophilic. For a deeper understanding of these metrics, we suggest the interested reader refer to the original paper [Platonov et al., 2023].
Datasets $N$ $|\mathcal{E}|$ $d$ $|C|$ $\xi_{edge}$ $\xi_{adj}$
Amazon Ratings 24492 186100 300 5 0.38 0.14
Roman Empire 22662 65854 300 18 0.05 -0.05
Minesweepers 10000 78804 7 2 0.68 0.01
Questions 48921 307080 301 2 0.84 0.02
Tolokers 11758 1038000 10 2 0.59 0.09
Table 5: Dataset Information
The dataset released by [Dwivedi et al., 2023] instead has the statistics reported in Tables 6 and 7.
Table 6: Dataset details, tasks, and performance metrics.
Dataset Domain Task Node Features (dim) Edge Features (dim) Performance Metric
PascalVOC-SP Computer Vision Node Classification Pixel + Coord (14) Edge Weight (1 or 2) macro F1
Peptides-func Chemistry Graph Classification Atom Encoder (9) Bond Encoder (3) AP
Peptides-struct Chemistry Graph Regression Atom Encoder (9) Bond Encoder (3) MAE
Table 7: Statistics of the proposed LRGB datasets.
Dataset Total Graphs Total Nodes Avg Nodes Mean Deg. Total Edges Avg Edges Avg Shortest Path Avg Diameter
PascalVOC-SP 11,355 5,443,545 479.40 5.65 30,777,444 2,710.48 $10.74\pm 0.51$ $27.62\pm 2.13$
Peptides-func 15,535 2,344,859 150.94 2.04 4,773,974 307.30 $20.89\pm 9.79$ $56.99\pm 28.72$
Peptides-struct 15,535 2,344,859 150.94 2.04 4,773,974 307.30 $20.89\pm 9.79$ $56.99\pm 28.72$
E.2 Baselines
Baselines for heterophilic graphs. We consider the set of heterophilic graphs introduced in Platonov et al. [2023]. We test our models in heterophilic settings, since these are scenarios that typically cause over-smoothing, and we want to gauge in practice how they compare against existing models. We briefly describe the baselines used for comparison against SAS-GNN and EEGNN in node classification. From the benchmark introduced in [Platonov et al., 2023], we include the reported performances of GCN, GraphSAGE [Kipf and Welling, 2017, Hamilton et al., 2017], GAT [Veličković et al., 2018], and GT [Dwivedi and Bresson, 2021a]. Additionally, we consider GAT-sep and GT-sep, two variants of these models specifically designed for heterophily. Even though these baselines seem outdated, it is known that the experiments in [Platonov et al., 2023] revealed that classic models outperform methods specialized for heterophily. More recently, Finkelshtein et al. [2024] introduced the Cooperative Graph Neural Networks (Co-GNNs) model, which we also include in our evaluation. This family of models is also more expressive than classic message-passing neural networks [Xu et al., 2019], thanks to an asynchronous message-passing.
Baselines for Long Range Graph Benchmark. We compare SAS-GNN and EEGNN on 3 datasets from LRGB Dwivedi et al. [2023]. These datasets were collected to assess how much GNNs can preserve information exchanged among distant nodes. We take advantage of these datasets to test how much our model performs compared to other methods. Among the baselines, we compare Classic MPNNs: GCN Kipf and Welling [2017], GIN Xu et al. [2019], GINE Hu et al. [2020], GatedGCN Bresson and Laurent [2018], which is the category that encompasses SAS-GNN and EEGNN. These methods need to process $l$ message-passing steps to allow nodes distant $l$ to communicate. Graph Transformers: Dwivedi and Bresson [2021a], SAN Kreuzer et al. [2021], and GraphGPS Rampášek et al. [2023], this category uses self-attention to all the nodes, so it allows nodes to reach every part of the graph, but it has a quadratic complexity w.r.t $|\mathcal{V}|$. Rewiring methods: DIGL Gasteiger et al. [2022], MixHop Abu-El-Haija et al. [2019], and DRew Gutteridge et al. [2023], these approaches modify the graph topology during the message-passing, in this way messages can be delivered through paths that did not exist before, and so long-range interactions are fostered. Asynchronous Message-Passing methods: Co-GNN Finkelshtein et al. [2024], AMP Errica et al. [2023]. These can be seen similarly to graph rewiring, since the computational graph is different from the input graph, as well as rewiring methods. However, here there exists the interpretation of nodes that decides what messages to filter, rather than rewiring the graph.
We test our approach only on 3 out of 5 datasets, because of computational constraints. These are Peptides-func, Peptides-struct, and Pascal-VOC.
E.3 Implementation Details
The metrics that we use are the Area Under the Curve (AUC) for the binary classification experiments, accuracy for Roman Empire and Amazon Ratings, average precision (AP) for Peptides-func, mean-absolute error (MAE) for Peptides-struct, and macro F1-Score (F1) for Pascal-VOC. All the experiments concerning EEGNN and SAS-GNN implied that we removed self-loops from the graphs, in order to satisfy the assumptions made in Theorems 3.1, 3.2. We always apply $f(\mathbf{X})$ implemented via a one-layer perception with ReLU. And in the LRGB, we also use a final decoder MLP to refine the representations of $\mathbf{Z}$, details on the hyperparameter are in the code. In the heterophilic graphs, each training lasts 3000 epochs, and the optimizer that we use is Adam [Kingma and Ba, 2017]. Concerning the quantitative results, we report the mean and standard deviation, which are computed from the metric averaged across 10 runs. Each run corresponds to a different split of the nodes used for training, validation, and testing. These experiments were run on a single Nvidia GeForce RTX 3090 Ti 24 GB and also a single NVIDIA A100 80GB PCIe. As concerns the LRGB experiments, we run each configuration for a maximum of 1000 epochs across 4 different weight initialization, using the AdamW optimizer. We picked the best configuration based on the validation metric. These experiments were run on a single NVIDIA A100 80GB PCIe. As concerns with the experiments on the TUDatasets for short-range graph classification (see Table 17) and homophilic node classification (see Section F.7), we also used a single NVIDIA A100 80GB PCIe. All the data was taken mostly from the PyTorch Geometric library [Fey and Lenssen, 2019], and other open-source sources. We always refer to their provenance in the code when using them.
E.4 Hyperparameters
For a fair comparison, the hyperparameter tuning process was done according to the hyperparameter set used by Finkelshtein et al. [2024]. In the SAS-GNN experiments, we choose $\tau$ among $\{0.1,0.5,1\}$. While in the experiments with EEGNNs, we do not need to tune $\tau$ because it is adaptive (see Algorithm 1), and we do not need to tune the number of layers parameter since we use a large and upper bound depth $L=20$. In the case of EEGNN, we choose $f_{c}(\cdot)$ as a Mean-GNN [Hamilton, ]. Similarly to Finkelshtein et al. [2024], we test a contained number of layers $\{1,2,3\}$, and then as hidden dimension we select $4$, $8$, $16$, $32$ or $64$. Then we also tune the value of the temperature $\nu_{0}$ between 00 and 0.1. $f_{\nu}(\cdot)$, the network that learns the adaptive temperature $\nu$, adopts the design proposed in [Finkelshtein et al., 2024]. In the experiments concerning the LRGB datasets, $f_{c}$ and $f_{\nu}$ keep the same hyperparameter set as the heterophilic experiments. The main hyperparameter set is the same used in Tönshoff et al. [2023]. In any experiment with SAS-GNN and EEGNN, we do not use any dropout or normalization layer within the message-passing updates, to follow the ODE interpretation and directly minimize the energy functional in Equation (3). We only use them in the feature encoder and in the final decoder. In the experiments with Peptides-func and Peptides-struct, we use two positional encoding techniques, which are typically approached in these settings. For Peptides-func we used RWSE [Dwivedi et al., 2022], while instead in Peptides-struct, we used LapPE [Dwivedi and Bresson, 2021b]. We used these according to the best hyperparameters used for graph convolutional methods in Tönshoff et al. [2023]. We do not use positional encodings in Pascal-VOC.
Appendix F Additional Results
F.1 SAS-GNN is able to retain performance as depth increases
Having proposed SAS-GNN, we have demonstrated that it is always stable and non-dissipative, and also able to allow each node to induce attraction or repulsion edge-wise. We study the behavior of SAS-GNN when the depth increases up to 50 layers, which is an overestimate of what we need in practice. In our experiments, we design SAS-GNN, as specified in Corollary 3.3. All the data points corresponds to a different training procedure, where we display the best accuracy. We present the results in Figure 4(a), where we show an experiment with Cora, a homophilic dataset. In Figure 4(b), we assess the resilience to depth via Minesweeper, an example of a heterophilic graph. In Cora, we notice that both A-DGN and SAS-GNN can retain their performance. Also, GRAFF can maintain it up to the 30th layer, then it starts decreasing. This behavior is expected since GRAFF is only designed to contrast over-smoothing, which does not occur, but it is not specifically designed to be robust to the number of layers. It is interesting to notice that SAS-GNN, even though part of its architecture is inherited by GRAFF, can retain its performance, behaving similarly to A-DGN. As far as GCN is concerned, we see that it has started losing its performance in the earlier layers. Here, we consider a GCN design that lacks the residual connection and does not use weight sharing. Furthermore, Table 8 presents an individual analysis of SAS-GNN on the Minesweeper dataset as the number of layers increases up to 100. In this experiment, we also allow the hidden dimension $m^{\prime}$ to scale to assess the robustness of the model under varying capacity. We observe that the model maintains its performance as the number of layers increases. However, increasing $m^{\prime}$ from 32 to 256 leads to a drop in performance. We attribute this decline to reduced generalization at higher hidden dimensions. Notably, $m^{\prime}=32$ emerged as the optimal setting based on our hyperparameter tuning procedure.
Table 8: Performance across different model sizes and layer depths
Model 10 layers 50 layers 100 layers # of Parameters
$m^{\prime}=32$ 92.95 $\pm$ 0.5 93.05 $\pm$ 0.3 93.64 $\pm$ 0.2 2,496
$m^{\prime}=64$ 91.95 $\pm$ 0.1 92.54 $\pm$ 0.1 92.85 $\pm$ 0.3 9,088
$m^{\prime}=128$ 91.20 $\pm$ 0.2 92.87 $\pm$ 0.1 93.24 $\pm$ 0.1 34,560
$m^{\prime}=256$ 90.73 $\pm$ 0.1 92.67 $\pm$ 0.2 92.90 $\pm$ 0.3 134,656
Figure 4: Comparison of models’ performance on different datasets with increasing depth.
(a) Cora: Comparison of models’ performance w.r.t increasing depth, in a homophilic graph.
(b) Minesweeper: Comparison of models’ performance w.r.t increasing depth under heterophily.
F.2 Extended experimental proofs on Over-smoothing and Over-Squashing mitigation
Here, we provide additional plots of Dirichlet Energy and layer-wise sensitivity to witness the model’s resilience to over-smoothing and over-squashing. Here we perform one training per model, sharing for each model the hyperparameters in Table 9. GCN is implemented such that over-smoothing is prone to occur, so we did not use residual connections and used ReLU as activation function [Scholkemper et al., 2024, Álvaro Arroyo et al., 2025].
Table 9: Hypeparameters used for the Dirichlet Energy and Sensitivity analyses.
Amazon Ratings Roman empire Minesweeper Questions Tolokers Peptides-func Peptides-struct Pascal-VOC
$m^{\prime}$ 256 128 32 64 32 300 300 128
$L$ 50 50 50 50 50 10 10 10
We report the sensitivity $S$ for Pascal-VOC and Peptides-struct in Figure 5. We notice that, similarly to the example in the main paper, we do not record any difficulties by the models to keep high sensitivity. Also, we notice that in Peptides-struct, we have a similar trend to Peptides-func, which reflects the same dataset domain. In Figure 5(a), we record a decreasing trend in sensitivity from all the models, but in general, they all have $S_{l}\neq 0$, also in this case. In Figure 6, we analyse the $E^{dir}$ trends for GCN, GRAFF, A-DGN, and SAS-GNN, to understand whether SAS-GNN effectively mitigates over-smoothing. We observe that also in this case, as in the Minesweeper example of Figure 2(a), SAS-GNN follows a similar trend of A-DGN, in all the other datasets, meaning that its ability to mitigate over-smoothing is more inherited from this model, rather than GRAFF. In these other experiments, we notice that $E^{dir}$ for GRAFF grows exponentially. As observed also in Minesweeper, it is clear that these plots describe the over-sharpening phenomenon discussed by Di Giovanni et al. [2023]. For instance, GRAFF is a more expressive version of GCN, which is provably capable of both asymptotically experiencing over-smoothing, as well as the opposite effect, namely over-sharpening. The desired behavior is a trade-off of the two, which is what we observe when GRAFF has a contained number of layers. Here, GRAFF underperforms due to an asymptotic behavior that lets adjacent node features diverge; indeed, $E^{dir}$ increases. As concerns GCN, as expected, it experiences over-smoothing also in the remaining heterophilic datasets.
(a) Pascal-VOC: Sensitivity analysis in node classification on LRGB dataset.
(b) Peptides-struct: Sensitivity analysis in node classification on LRGB dataset.
Figure 5: Comparison of models’ performance on different datasets with increasing depth.
(a) $E^{dir}$ trend for Tolokers.
(b) $E^{dir}$ trend for Questions.
(c) $E^{dir}$ trend for Roman Empire.
(d) $E^{dir}$ trend for Amazon Ratings.
Figure 6: $E^{dir}$ trends for heterophilic graphs, with associated accuracies.
F.3 Analysis on the use of ReLU+TanH as activation
Here, we want to understand how our instantiation of the non-linearities $\sigma_{1}(x)=\text{ReLU}(\text{TanH}(x))$ and $\sigma_{2}=\text{ReLU}(x)$ results in effective w.r.t. more common choices in the design of neural networks, such as ReLU or TanH. In order to assess these aspects, we compare SAS-GNN against A-DGN and GRAFF, using all the possible combinations of ReLU, TanH, and ReLU+TanH. Our objective is to understand whether we have any advantage in terms of performance, rather than a result that is limited to the theory. We present such a comparison in Table 10. From the Table, we see that our principled approach ranks first in 2 out of 5 datasets, namely Tolokers and Questions. While A-DGN performs significantly better than GRAFF and SAS-GNN in Minesweeper and Roman Empire. In the other datasets, the performance remains comparable for all the activation types. Surprisingly, GRAFF tends to perform worse than the other models, even though it was designed specifically to handle heterophilic graphs. However, when it was proposed by Di Giovanni et al. [2023], this specific benchmark was not introduced yet, thus, we attribute this gap to this new collection. We present additional evidence on the impact of ReLU+TanH, when used within other GNNs, such as GCN, GraphSAGE, GAT, and GIN, in Section F.6.
Table 10: Comparison of the activation types based on the models. The scores are marked in red for the first, blue for the second. These results are taken from Platonov et al. [2023], Finkelshtein et al. [2024], Luo et al. [2024].
Model Activation Amazon Ratings Minesweeper Roman Empire Tolokers Questions
A-DGN TanH 49.14 $\pm$ 0.7 94.81 $\pm$ 0.4 82.53 $\pm$ 0.7 84.90 $\pm$ 1.1 76.51 $\pm$ 1.0
ReLU 51.41 $\pm$ 0.4 95.08 $\pm$ 0.5 84.83 $\pm$ 0.6 85.32 $\pm$ 1.0 78.24 $\pm$ 1.8
ReLU+TanH 51.57 $\pm$ 1.3 94.86 $\pm$ 0.4 84.66 $\pm$ 0.6 85.35 $\pm$ 1.7 79.01 $\pm$ 1.1
GRAFF TanH 50.02 $\pm$ 1.5 92.44 $\pm$ 0.7 78.76 $\pm$ 0.8 85.03 $\pm$ 1.8 76.77 $\pm$ 1.5
ReLU 51.41 $\pm$ 0.7 93.23 $\pm$ 0.5 78.98 $\pm$ 0.4 85.67 $\pm$ 0.7 76.27 $\pm$ 1.2
ReLU+TanH 50.77 $\pm$ 0.6 92.59 $\pm$ 0.7 78.59 $\pm$ 0.5 85.29 $\pm$ 0.9 79.04 $\pm$ 1.1
SAS-GNN TanH 48.91 $\pm$ 0.8 93.80 $\pm$ 0.5 82.07 $\pm$ 0.5 84.56 $\pm$ 0.7 77.93 $\pm$ 1.3
ReLU 51.39 $\pm$ 0.6 93.38 $\pm$ 0.5 83.59 $\pm$ 0.5 85.72 $\pm$ 1.8 79.13 $\pm$ 0.4
ReLU+TanH 51.47 $\pm$ 0.7 93.29 $\pm$ 0.6 83.46 ± 0.6 85.80 $\pm$ 0.8 79.60 $\pm$ 1.1
F.4 Full Results on the benchmark for heterophily
In this section, we present the full leaderboard results from the benchmark proposed by Platonov et al. [2023]. While Table 3 already reported partial results showcasing the best-performing models, here we provide a more complete view.
The benchmark from Platonov et al. [2023] focuses on evaluating GNNs on a suite of diverse, medium-to-large graphs with varying degrees of heterophily, domain, density, and size. In their study, it was demonstrated that many GNN models originally designed to handle heterophily underperform compared to classical GNNs, such as GCN, GraphSAGE, and GAT, or even GTs, when evaluated under this more realistic setting.
Given this context, our proposed SAS-GNN and EEGNN models exhibit superior performance relative to heterophily-specialized methods, and are comparable to, or even outperform, the classic GNNs in several cases. Additionally, we include results for the same classical models after undergoing a more extensive hyperparameter tuning, as reported by Luo et al. [2024]. These tuned versions incorporate techniques such as dropout, batch normalization, and residual connections, which lead to significant performance improvements for both message-passing architectures (e.g., GCN, GraphSAGE, GAT) and Graph Transformers (e.g., Polynormer).
Despite not employing normalization or dropout in our models, SAS-GNN and EEGNN remain competitive on Tolokers and Questions, where we still match or surpass several strong baselines. However, on other datasets, we acknowledge a noticeable gap when compared to extensively tuned models. This observation further highlights the importance of architectural principles that incorporate mechanisms like weight sharing and depth-resilient message-passing. This enables our models to maintain competitive performance in diverse settings without relying on auxiliary components.
Table 11: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. The models marked with the asterisks are those presented in Table 3, reported by Luo et al. [2024].
Model Amazon Ratings Minesweeper Roman Empire Tolokers Questions
Classic GNNs Platonov et al. [2023]
GCN 48.70 ± 0.63 89.75 ± 0.52 73.69 ± 0.74 83.64 ± 0.67 76.09 ± 1.27
SAGE 53.63 ± 0.39 93.51 ± 0.57 85.74 ± 0.67 82.43 ± 0.44 76.44 ± 0.62
GAT 49.09 ± 0.63 92.01 ± 0.68 80.87 ± 0.30 83.70 ± 0.47 77.43 ± 1.20
GAT-sep 52.70 ± 0.62 93.91 ± 0.35 88.75 ± 0.41 83.78 ± 0.43 76.79 ± 0.71
GT 51.17 ± 0.66 91.85 ± 0.76 86.51 ± 0.73 83.23 ± 0.64 77.95 ± 0.68
GT-sep 52.18 ± 0.80 92.29 ± 0.47 87.32 ± 0.39 82.52 ± 0.92 78.05 ± 0.93
Heterophily-Specialized Models Platonov et al. [2023]
H2GCN 36.47 ± 0.23 89.71 ± 0.31 60.11 ± 0.52 73.35 ± 1.01 63.59 ± 1.46
CPGNN 39.79 ± 0.77 52.03 ± 5.46 63.96 ± 0.62 73.36 ± 1.01 65.96 ± 1.95
GPR-GNN 44.88 ± 0.34 86.24 ± 0.61 64.85 ± 0.27 72.94 ± 0.97 55.48 ± 0.91
FSGNN 52.74 ± 0.83 90.08 ± 0.70 79.92 ± 0.56 82.76 ± 0.61 78.86 ± 0.92
GloGNN 36.89 ± 0.14 51.08 ± 1.23 59.63 ± 0.69 73.39 ± 1.17 65.74 ± 1.19
FAGCN 44.12 ± 0.30 88.17 ± 0.73 65.22 ± 0.56 77.75 ± 1.05 77.24 ± 1.26
GBK-GNN 45.98 ± 0.71 90.85 ± 0.58 74.57 ± 0.47 81.01 ± 0.67 74.47 ± 0.86
JacobiConv 43.55 ± 0.48 89.66 ± 0.40 71.14 ± 0.42 68.66 ± 0.65 73.88 ± 1.16
Reassessment + Current SOTA Luo et al. [2024]
GCN* 53.80 ± 0.60 97.86 ± 0.52 91.27 ± 0.20 83.64 ± 0.67 79.02 ± 1.27
SAGE* 55.40 ± 0.21 97.77 ± 0.62 91.06 ± 0.67 82.43 ± 0.44 77.21 ± 1.28
GAT* 55.54 ± 0.51 97.73 ± 0.73 90.63 ± 0.14 83.78 ± 0.43 77.95 ± 0.51
GT 51.17 ± 0.66 91.85 ± 0.76 86.51 ± 0.73 83.23 ± 0.64 77.95 ± 0.68
Polynormer* 54.81 ± 0.49 97.46 ± 0.36 92.55 ± 0.37 85.91 ± 0.74 78.92 ± 0.89
CO-GNN 54.17 ± 0.37 97.31 ± 0.41 91.37 ± 0.35 84.45 ± 1.17 76.54 ± 0.95
Ours Luo et al. [2024]
SAS-GNNnoedge 51.47 ± 0.68 93.29 ± 0.61 83.46 ± 0.61 85.80 ± 0.79 79.60 ± 1.15
EEGNN 51.47 ± 0.51 93.18 ± 1.37 80.36 ± 0.43 85.26 ± 0.65 78.90 ± 1.15
F.5 Full Results on Long Range Graph Benchmark
In the main paper, we discussed the performance of SAS-GNN and EEGNN against MPNNs and GTs, as well as asynchronous message-passing methods such as Co-GNN and AMP. We limited our analysis, since we consider them as our main competitor on the LRGB task. However, other approaches exist that attempt to improve the GNNs’ capabilities to overcome their long-range information propagation capabilities. Among these, we have Rewiring Methods that also contribute to high scoring in long-range tasks. Here we include DIGL [Gasteiger et al., 2022], MixHop [Abu-El-Haija et al., 2019], and DRew [Gutteridge et al., 2023]. The results are in Table 12. We also distinguish here between SAS-GNNnoedge ($f_{e}(\mathbf{E})=0$), SAS-GNNedge ($f_{e}(\mathbf{E})=\mathbf{BEW}_{e}$), and SAS-GNN/EEGNNours ($f_{e}(\mathbf{E})=-\text{ReLU}(\mathbf{BEW}_{e})$). While we do not surpass all the models, we outperform classic MPNNs in Peptides-func and Peptides-struct with SAS-GNNnoedge. Notably, edge features do not enhance performance in peptide datasets, though our edge-based method slightly improves over SAS-GNNedge. However, in Pascal-VOC, classic MPNNs achieve better results, yet our edge feature strategy improves F1, highlighting the informativeness of edge features in this domain. Against rewiring methods, we remain competitive. In Peptides-func, we surpass most models except DRew-GCN and its LapPE variant. In Peptides-struct, SAS-GNNedge underperforms, but all other versions achieve superior results. While DRew-GCN excels in peptide tasks, we outperform it in Pascal-VOC. Compared to DIGL, we perform better in peptide datasets but fall behind in Pascal-VOC. As concerns GTs, we outperform them in peptide settings except in inductive node classification. Against asynchronous MPNNs, we are slightly worse than Co-GNN in Peptides-func and comparable to AMP in Peptides-struct, despite these models being more expressive than the 1-WL test. No comparisons are available for Pascal-VOC in this category.
Table 12: Performance comparison of models across LRGB datasets. The scores are marked in red for the first, blue for the second, and green for the third.
Model Peptides-func Peptides-struct Pascal VOC-SP
AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$
Classic MPNNs
GCN 68.60±0.50 0.2460±0.0013 20.78±0.31
GINE 66.21±0.67 0.2473±0.0017 27.18±0.54
GatedGCN 67.65±0.47 0.2477±0.0009 38.80±0.40
Rewiring Methods
DIGL+MPNN 64.69±0.19 0.3173±0.0007 28.24±0.39
DIGL+MPNN+LapPE 68.30±0.26 0.2616±0.0018 29.21±0.38
MixHop-GCN 65.92±0.36 0.2921±0.0023 25.06±1.33
MixHop-GCN+LapPE 68.43±0.49 0.2614±0.0023 22.18±1.74
DRew-GCN 69.96±0.76 0.2781±0.0028 18.48±1.07
DRew-GCN+LapPE 71.50±0.44 0.2536±0.0015 18.51±0.92
GTs
GT+LapPE 63.26±1.26 0.2529±0.0016 26.94±0.98
SAN+LapPE 63.84±1.21 0.2683±0.0043 32.30±0.39
GraphGPS+LapPE 65.35±0.41 0.2500±0.0005 37.48±1.09
Asynchronous MPNNs
CO-GNNs 69.90 ± 0.93 - -
AMP 71.63 ± 0.58 0.2431 ± 0.0004 -
Ours
SAS-GNNnoedge 69.71 ± 0.62 0.2449 ± 0.0013 22.65 ± 0.27
SAS-GNNedge 69.27 ± 0.58 0.2547 ± 0.0163 23.31 ± 0.49
SAS-GNNours 69.44 ± 0.63 0.2528 ± 0.0130 25.64 ± 0.63
EEGNNours 68.23 ± 0.37 0.2532 ± 0.0050 24.10 ± 0.73
F.6 Results using Straight Trough Gumbel Softmax on Classic GNNs
In this section, we demonstrate that our early-exit module, based on the Straight-Through Gumbel-Softmax Estimator, is modular and can be integrated with various GNN architectures, including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], GAT [Veličković et al., 2018], and GIN [Xu et al., 2019].
[…]
All models use the same hyperparameters as the best-performing EEGNN configuration in Table 3; these configurations can be found in the code repository. It is important to note that these implementations differ from those in prior benchmarks such as Luo et al. [2024], as we apply weight sharing and operate within our specific experimental framework.
Our goal is to analyze how integrating the early-exit (EE) module affects each GNN’s performance. Table 13 summarizes this comparison. We highlight in bold all instances where using the EE module leads to performance improvements. These gains can be found more with models trained with ReLU, and are particularly pronounced when the base model struggles (e.g., GAT on Minesweeper or GIN on Questions). We hypothesize that these improvements stem from avoiding excessive depth, which otherwise leads to performance degradation due to over-processing. In such cases, the EE module enables early termination and stabilizes performance.
Notably, SAS-GNN does not suffer from such degradation, as shown earlier. Its design inherently mitigates over-smoothing and over-squashing, making it robust to deeper configurations. Conversely, the EE module can negatively impact performance in some setups—for instance, GIN on Minesweeper and GraphSAGE on Tolokers. This degradation likely arises from applying weight sharing to models not originally designed for it, potentially reducing their expressivity due to fewer parameters. Still, exceptions exist—e.g., GraphSAGE achieves unexpectedly high AUROC on Minesweeper, comparable to the values in Table 3.
Since SAS-GNN and EEGNN benefit from using a combination of ReLU and Tanh nonlinearities, as motivated in Corollary 3.3, we extended this configuration to the other GNNs. In Table 13, we mark with an asterisk all cases where ReLU+Tanh improves performance over ReLU alone. We find that this hybrid activation consistently enhances results across most settings.
To summarize, the early-exit module can significantly improve performance, particularly when the base model struggles with deeper architectures. In such cases, EE enables earlier termination and more stable outcomes. In other cases, both the base and EE-enhanced models may underperform, likely due to the limitations of weight sharing in architectures not originally designed for it. Nevertheless, we find evidence—such as GraphSAGE on Minesweeper—that weight sharing can still be effective on certain datasets. Additionally, using ReLU+Tanh, as proposed in SAS-GNN, proves beneficial even for classical GNNs, improving performance and narrowing the gap between base and EE-enhanced models.
Table 13: Performance of classic GNNs using Gumbel Softmax as early-exit. Models are also tested by changing the activation function type.
Model Minesweeper Questions Tolokers
ReLU ReLU + Tanh ReLU ReLU + Tanh ReLU ReLU + Tanh
GCN $91.29\pm 0.5$ $90.42\pm 0.6$ $74.92\pm 1.4$ ${}^{*}76.56\pm 1.0$ $75.58\pm 3.7$ ${}^{*}84.89\pm 0.89$
EE+GCN $91.07\pm 0.7$ $90.39\pm 0.6$ $\mathbf{75.56\pm 1.1}$ ${}^{*}76.33\pm 1.3$ $\mathbf{78.02\pm 5.3}$ ${}^{*}83.54\pm 1.4$
SAGE $95.61\pm 0.6$ ${}^{*}96.60\pm 0.3$ $75.03\pm 0.4$ ${}^{*}75.22\pm 1.0$ $84.54\pm 0.65$ $84.13\pm 0.42$
EE+SAGE $\mathbf{95.73\pm 0.4}$ ${}^{*}96.56\pm 0.6$ $71.36\pm 1.2$ ${}^{*}72.09\pm 1.7$ $76.04\pm 3.7$ $74.66\pm 0.99$
GAT $73.98\pm 15$ ${}^{*}91.08\pm 0.5$ $74.39\pm 0.9$ ${}^{*}75.53\pm 1.3$ $76.84\pm 7.8$ ${}^{*}84.09\pm 0.89$
EE+GAT $\mathbf{84.17\pm 14}$ ${}^{*}\mathbf{91.89\pm 1.5}$ $\mathbf{74.91\pm 1.1}$ $74.90\pm 2.2$ $\mathbf{79.02\pm 3.8}$ $80.44\pm 4.0$
GIN $61.92\pm 1.9$ ${}^{*}89.89\pm 0.7$ $50.0\pm 0.0$ ${}^{*}77.94\pm 1.2$ $50.0\pm 0.0$ ${}^{*}82.55\pm 1.1$
EE+GIN $51.11\pm 0.3$ $51.03\pm 1.6$ $\mathbf{71.27\pm 0.8}$ $70.99\pm 1.4$ $50.0\pm 0.0$ ${}^{*}78.87\pm 4.3$
SAS-GNN - $93.29\pm 0.61$ $79.60\pm 1.15$ $85.80\pm 0.79$
EEGNN - $93.18\pm 1.37$ $78.90\pm 1.15$ $85.26\pm 0.65$
F.7 Results on Homophilic Node Classification
To ensure completeness in terms of the graph types used for testing, we evaluate our models on standard homophilic node classification datasets. Our goal is to assess whether our models remain competitive, achieving performance comparable to classic GNNs. These benchmarks typically do not require deep architectures to achieve high accuracy, so we do not expect our methods to provide a clear advantage over shallow baselines in this setting.
We evaluate our proposed models, SAS-GNN and EEGNN, on transductive homophilic node classification on the datasets Computer, CS, Photo, and Physics, introduced by [Shchur et al., 2019]. We report the mean accuracy and standard deviation over 10 trials, with dataset statistics summarized in Table 14.
The baselines that we use are both message-passing methods, such as GCN, GraphSAGE, and GAT, as well as Graph Transformers (GT), such as NAGphormer Chen et al. [2023], Exphormer Shirzad et al. [2023], GOAT Kong et al. [2023], Polynormer Deng et al. [2024], and SGFormer Wu et al. [2024].
We used the same hyperparameter search space proposed by Luo et al. [2024]. Specifically, their configurations heavily rely on the use of normalization and dropout layers within the message-passing pipeline—techniques we deliberately avoid in favor of a principled design that emphasizes structural simplicity.
In Table 15, we show our models’ performance. We observe that they are never among the top-2 models. This is not surprising, since the other models in the benchmark leverage normalization and dropout layers, while our design is focused on efficiently handling long-range information propagation. Nevertheless, both SAS-GNN and EEGNN perform competitively, trailing the top models only marginally and outperforming several Transformer-based and classic baselines. These results further support the adaptability and robustness of our methods across diverse graph learning settings.
Table 14: Statistics of the homophlic node classification datasets.
Computer Photo CS Physics
# nodes 13,752 7,650 18,333 34,493
# edges 245,861 119,081 81,894 247,962
# features 767 745 6,805 8,415
# classes 10 8 15 5
metric Accuracy Accuracy Accuracy Accuracy
Table 15: Performance on Node classification datasets. These scores are taken from Luo et al. [2024]. The scores are marked in red for the first, blue for the second.
Computer Photo CS Physics
NAGphormer $91.69\pm 0.30$ $96.14\pm 0.16$ $95.85\pm 0.16$ $97.35\pm 0.12$
Exphormer $91.47\pm 0.17$ $95.35\pm 0.22$ $94.93\pm 0.01$ $96.89\pm 0.09$
GOAT $92.29\pm 0.37$ $94.33\pm 0.21$ $93.81\pm 0.19$ $96.47\pm 0.16$
GraphGPS $91.79\pm 0.63$ $94.89\pm 0.14$ $94.04\pm 0.21$ $96.71\pm 0.15$
Polynormer $93.78\pm 0.10$ $96.57\pm 0.23$ $95.42\pm 0.19$ $97.18\pm 0.11$
SGFormer $91.99\pm 0.76$ $95.10\pm 0.47$ $94.78\pm 0.20$ $96.60\pm 0.18$
GCN $93.99\pm 0.12$ $96.10\pm 0.46$ $96.17\pm 0.06$ $97.46\pm 0.10$
GraphSAGE $93.25\pm 0.14$ $96.78\pm 0.23$ $96.38\pm 0.11$ $97.19\pm 0.05$
GAT $94.09\pm 0.37$ $96.60\pm 0.33$ $96.21\pm 0.14$ $97.25\pm 0.06$
SAS-GNN $92.27\pm 0.29$ $96.47\pm 0.33$ $94.46\pm 0.10$ $96.49\pm 0.10$
EEGNN $90.24\pm 0.93$ $95.23\pm 0.35$ $95.56\pm 0.07$ $96.22\pm 0.09$
F.8 Results on TUDataset benchmark for Graph Classification
To ensure completeness, we evaluate our models on standard graph classification datasets. Our goal is to assess whether our models remain competitive, achieving performance comparable to classic GNNs. These benchmarks typically do not require deep architectures to achieve high accuracy, so we do not expect our methods to provide a clear advantage over shallow baselines in this setting.
We evaluate our proposed models, SAS-GNN and EEGNN, on the TUDataset graph classification benchmark [Morris et al., 2020], following the protocol established by Finkelshtein et al. [2024]. We report the mean accuracy and standard deviation across seven datasets, with dataset statistics summarized in Table 16.
Our models are compared against a broad range of baselines, including established GNNs such as DGCNN [Wang et al., 2019], DiffPool [Ying et al., 2019], ECC [Simonovsky and Komodakis, 2017], CGMM [Bacciu et al., 2020], ICGMMf [Castellana et al., 2022], SPN [Abboud et al., 2023], and GSPN [Errica and Niepert, 2024]. We also include results for Co-GNN [Finkelshtein et al., 2024] for direct comparison.
To ensure fairness, we adopt the same hyperparameter search space and evaluation pipeline used in previous work. Configurations that failed due to excessive memory or runtime demands are marked as OOR (Out of Resources) in the results tables.
As shown in Table 17, both SAS-GNN and EEGNN achieve competitive or superior performance across datasets. For instance, our models match or outperform Co-GNN on datasets like IMDB-B and PROTEINS. Notably, despite using simple mean pooling, our models sometimes surpass approaches with more complex pooling strategies such as DiffPool. Furthermore, EEGNN improves upon SAS-GNN on the ENZYMES dataset, despite the high variance typically observed there. The exception is NCI1, where exiting before the full-depth lead to a significant performance degradation.
Table 16: Dataset statistics for the TUDataset benchmark for graph classification.
IMDB-B IMDB-M REDDIT-B REDDIT-M NCI1 PROTEINS ENZYMES
# graphs 1000 1500 2000 4999 4110 1113 600
# avg. nodes 19.77 13.00 429.63 508.51 29.87 39.06 32.63
# avg. edges 96.53 65.94 497.75 1189.74 32.30 72.82 64.14
# classes 2 3 2 5 2 2 6
metrics Accuracy Accuracy Accuracy Accuracy Accuracy Accuracy Accuracy
Table 17: Graph classification results on the TUDataset benchmark. The scores are marked in red for the first, blue for the second.
IMDB-B IMDB-M REDDIT-B REDDIT-M NCI1 PROTEINS ENZYMES
DGCNN 69.2 $\pm$ 3.0 45.6 $\pm$ 3.4 87.8 $\pm$ 2.5 49.2 $\pm$ 1.2 76.4 $\pm$ 1.7 72.9 $\pm$ 3.5 38.9 $\pm$ 5.7
DiffPool 68.4 $\pm$ 3.3 45.6 $\pm$ 3.4 89.1 $\pm$ 1.6 53.8 $\pm$ 1.4 76.9 $\pm$ 1.9 73.7 $\pm$ 3.5 59.5 $\pm$ 5.6
ECC 67.7 $\pm$ 2.8 43.5 $\pm$ 3.1 OOR OOR 76.2 $\pm$ 1.4 72.3 $\pm$ 3.4 29.5 $\pm$ 8.2
GIN 71.2 $\pm$ 3.9 48.5 $\pm$ 3.3 89.9 $\pm$ 1.9 56.1 $\pm$ 1.7 80.0 $\pm$ 2.4 75.3 $\pm$ 2.5 59.6 $\pm$ 4.5
GraphSAGE 68.8 $\pm$ 4.5 47.6 $\pm$ 3.5 84.3 $\pm$ 1.9 50.0 $\pm$ 3.6 74.9 $\pm$ 1.9 70.5 $\pm$ 3.2 58.2 $\pm$ 6.0
CGMM - - 88.1 $\pm$ 1.9 - - - -
ICGMMf 71.8 $\pm$ 4.4 49.0 $\pm$ 3.8 91.6 $\pm$ 2.1 55.6 $\pm$ 1.7 76.4 $\pm$ 1.3 73.2 $\pm$ 3.9 -
SPN($k=5$) - - - - 74.2 $\pm$ 2.7 - 69.4 $\pm$ 6.2
GSPN - - 90.5 $\pm$ 1.1 55.3 $\pm$ 2.0 76.6 $\pm$ 1.9 - -
Co-GNN 72.2 $\pm$ 4.1 49.9 $\pm$ 4.5 90.5 $\pm$ 1.9 56.3 $\pm$ 2.1 79.4 $\pm$ 0.7 71.3 $\pm$ 2.0 68.3 $\pm$ 5.7
SAS-GNN 72.3 $\pm$ 2.9 48.0 $\pm$ 4.5 88.3 $\pm$ 2.3 55.1 $\pm$ 2.2 77.9 $\pm$ 1.8 72.0 $\pm$ 3.0 66.5 $\pm$ 5.9
EEGNN 71.2 $\pm$ 3.9 46.8 $\pm$ 4.4 86.5 $\pm$ 2.8 53.9 $\pm$ 2.4 62.53 $\pm$ 4.3 70.8 $\pm$ 2.7 70.3 $\pm$ 7.3
F.9 Extended Runtime Analysis
In Table 18, we show our extended runtime analysis, and in Table 19.
Table 18: Runtime Analysis Across Datasets and Layers.
Model Roman Empire Minesweeper Tolokers Questions
10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers
GCN $0.0266\pm 0.0108$ $0.0411\pm 0.0147$ $0.0139\pm 0.0082$ $0.0273\pm 0.0119$ $0.0286\pm 0.0099$ $0.0428\pm 0.0124$ $0.0168\pm 0.0012$ $0.0251\pm 0.0037$
Co-GNN $0.0553\pm 0.0167$ $0.0739\pm 0.0253$ $0.0315\pm 0.0091$ $0.0665\pm 0.0221$ $0.0498\pm 0.0156$ $0.0874\pm 0.0225$ $0.0352\pm 0.0038$ $0.0598\pm 0.0087$
Polynormer $0.0192\pm 0.0031$ $0.0312\pm 0.0045$ $0.01509\pm 0.0022$ $0.0274\pm 0.0030$ $0.0183\pm 0.0025$ $0.0322\pm 0.0045$ $0.0278\pm 0.0053$ $0.0381\pm 0.0078$
SAS-GNN $0.0323\pm 0.0126$ $0.0510\pm 0.0156$ $0.0177\pm 0.0080$ $0.0371\pm 0.0123$ $0.0268\pm 0.0098$ $0.0437\pm 0.0131$ $0.0198\pm 0.0023$ $0.0290\pm 0.0026$
EEGNN $0.0288\pm 0.0107$ $0.0290\pm 0.0117$ $0.0153\pm 0.0108$ $0.0209\pm 0.0111$ $0.0227\pm 0.0120$ $0.0283\pm 0.0126$ $0.0302\pm 0.0210$ $0.0327\pm 0.0263$
Model Roman Empire Minesweeper Tolokers Questions
10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers
GCN 20,768 31,328 10,880 21,440 10,976 21,536 20,288 30,848
Co-GNN 35,478 56,278 25,574 46,374 25,670 46,470 34,982 55,782
Polynormer 48,164 81,124 37,732 70,692 37,828 70,788 50,404 80,100
SAS-GNN 12,320 12,320 2,432 2,432 2,528 2,528 11,840 11,840
EEGNN 14,562 14,562 4,674 4,674 4,770 4,770 14,082 14,082
Table 19: Number of Parameters Across Datasets and Layers.
We confirm our claims in Section 4.1, where we show that not only are SAS-GNN and EEGNN parameter-efficient, but EEGNN preserves its time complexity even when allowing it to go deeper, implementing it with a higher $L$ budget (i.e., from $L=10$ to $L=20$). This analysis can be seen as a way to say that EEGNN has constant time complexity w.r.t. the number of layers. This is simply a result of the learning algorithm that does not let the nodes/graph exit distribution change when the budget has increased. However, whenever we record an increased inference time by EEGNN, it could mean that the nodes/graphs require could benefit from more processing time, and increasing $L$ allows the model to test deeper configurations.
F.10 Promising Directions for Early-Exit and GNNs
In addition to the results presented in the main paper, we aimed to explore whether implementing early-exit mechanisms in GNNs has space for accuracy improvements. Spinelli et al. [2021] provides evidence of accuracy enhancements within their experiments, but their experimental setup differs from ours; in particular, we consider heterophilic graphs and do not use any additional losses. However, due to the depth robustness of SAS-GNN, we hypothesize that exiting early or late may offer no significant accuracy gains, as the network’s performance is stable across varying depths. To assess this hypothesis, we designed an oracle-like evaluation for SAS-GNN. In this setup, the network ‘knows’ the optimal exit point for each node, defined as the point where its latent representation would be classified correctly. Our goal is to compare this oracle-based performance with that of a standard SAS-GNN. If the accuracy remains the same, we conclude that EEGNN does not provide any inherent accuracy advantage. More specifically, for each node, we track the logits produced at each possible exit point and check whether they yield a correct prediction. If no intermediate layer produces a correct output, we default to the logits from the final layer. The results of this analysis are summarized in Table 20.
Table 20: Metrics computed at the Optimal Exit.
Metrics No Exit Optimal Exit
Amazon Ratings 51.47 ± 0.68 68.36 ± 0.86
Tolokers 85.80 ± 0.79 89.21 ± 1.37
Questions 79.60 ± 1.15 80.35 ± 1.29
Roman Empire 83.46 ± 0.61 89.45 ± 0.37
Minesweeper 93.29 ± 0.61 96.93 ± 0.58
As seen in the Table, individual nodes often arrive at correct predictions earlier in the forward pass, but when forced to process through all layers, the overall performance drops compared to the optimal exit scenario. This aligns with the ‘over-thinking’ phenomenon frequently observed in early-exit neural networks [Wang et al., 2017], where intermediate layers produce accurate predictions, but later layers overwrite them with incorrect outputs. Now that GNNs have made progress in mitigating simple message-passing flaws and deepening architectures, these results highlight the potential of EEGNNs to further enhance performance by integrating early-exit strategies. We believe that this promising direction opens new possibilities for more efficient and accurate GNNs. So far, we observed that our current framework tends only to reconstruct the performance of SAS-GNN, but the results presented in Table 20 let us believe that EEGNN can impact more than what is currently achieved.
Table 21: Comparison of the fixed number of layers selected by SAS-GNN (i.e., Fixed # Layers) with the minimum, median, and maximum number of layers selected by EEGNN (Min. Layers, Median Layers, Max. Layers). The values are averaged across 10 test splits and are taken from the discrete exit point distributions.
Datasets Fixed # Layers Min. Layers Median Layers Max. Layers
Amazon Ratings 10 5.1 $\pm$ 2.64 13.8 $\pm$ 1.39 21.0 $\pm$ 0.00
Roman Empire 12 1.7 $\pm$ 0.48 3.3 $\pm$ 0.48 10.2 $\pm$ 4.84
Minesweeper 15 11.4 $\pm$ 10.13 15.6 $\pm$ 8.69 15.7 $\pm$ 8.53
Questions 9 1.0 $\pm$ 0.00 7.0 $\pm$ 0.66 18.8 $\pm$ 2.78
Tolokers 10 1.7 $\pm$ 0.82 8.3 $\pm$ 2.00 21.0 $\pm$ 0.00
Figure 7: Node Exit per Layer discrete distribution in Pascal-VOC.
Figure 8: Tolokers: Exit point of the nodes in the test set. EEGNNs let the nodes choose their exit in the continuous domain (Left). We can visualize their exit in the discrete domain as well (Right).
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 9: Amazon Ratings: Exit point of the nodes in the test set. We have the discrete case (Right) and the continuous case (Left), thanks to our neural Adaptive-step mechanism.
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 10: Minesweeper: Exit point of the nodes in the test set. We have the discrete case (Right) and the continuous case (Left), thanks to our neural Adaptive-step mechanism. At the first fold.
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 11: Minesweeper: Exit point of the nodes in the test set. We have the discrete case (Right) and the continuous case (Left), thanks to our neural Adaptive-step mechanism. At the second fold.
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 12: Minesweeper: Exit point of the nodes in the test set. We have the discrete case (Right), and the continuous case (Left), thanks to our neural Adaptive-step mechanism. At the fifth fold.
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 13: Questions: Exit point of the nodes in the test set. We have the discrete case (Right) and the continuous case (Left), thanks to our neural Adaptive-step mechanism.
(a) Node Exit in the time domain.
(b) Node exits at each layer.
Figure 14: Roman Empire: Exit point of the nodes in the test set. We have the discrete case (Right) and the continuous case (Left), thanks to our neural Adaptive-step mechanism.
F.11 Additional examples of exit distributions
In the main paper, we provide an example of the graph exit distribution in the test set from Pascal-VOC. Then we show these comparisons for Tolokers, Amazon Ratings, Minesweeper, Roman Empire, and Questions. We want to highlight 2 main aspects. 1) In most cases, nodes tend to exit before and after the fixed-depth selected in the vanilla SAS-GNN. 2) These distributions may differ from split to split. Since each node may require a different processing time, also each split will have different distributions as well.
We report the Pascal-VOC graphs exit distribution in Figure 7. In Table 21, we compare the fixed depth chosen in our SAS-GNN experiments with the discrete node distributions statistics, such as the minimum, maximum, and median exit point. We average these from each test split and report the mean and standard deviation. For EEGNN, we chose a fixed budget of 20 layers, which we see from our experiments was enough w.r.t. the median of the distribution. From Table 21, we have both cases where the median of the distributions is higher than the fixed depth, such as Amazon Ratings (see Figure 9), but also cases where the nodes tend to exit even before the fixed number of layers such as in Roman Empire and Questions, see Figures 14, and 13. As shown in Figure 8 for Tolokers, many nodes exit before $l=10$, while others continue beyond this threshold. In contrast, SAS-GNN enforces a fixed $L=10$, constraining the model to a predetermined behavior. In Appendix F.11, we further detail node exit distributions across datasets, showing varied exit patterns. The standard deviations in the table instead imply that these distributions vary across different test splits. In fact, according to our framework we expect all the nodes to behave differently, so also across the splits we expect different distributions. The Minesweeper results report a dramatically high standard deviation. If we analyze the distribution plots of Minesweeper, we understand the reason for this high standard deviation. In particular, across different splits, we see that in some cases all the nodes exit early in Figure 11, do not exit at all in Figure 10, or the exit distribution is smooth 12. Thanks to EEGNN, each node, based on its exit need, can cover this set of behaviors.
Appendix G Extended Related Work
This section reviews the literature most relevant to our work, encompassing Graph Neural Networks (GNNs), advanced message-passing strategies, early-exit mechanisms for GNNs, and early-exit techniques in general neural architectures.
Graph Neural Networks (GNNs). Most GNNs follow the message-passing paradigm [Gilmer et al., 2017], where node features are iteratively updated by aggregating information from neighboring nodes via convolutions, attention, or neural networks [Veličković, 2023]. These Message-Passing Neural Networks (MPNNs), including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], and GAT [Veličković et al., 2018], achieve strong empirical performance but are inherently limited by the number of layers: a $k$-layer MPNN can only propagate information along paths of length at most $k$. Deep MPNNs often suffer from over-smoothing (node features become indistinguishable) [Rusch et al., 2023, NT and Maehara, 2019], over-squashing (long-range information compressed through narrow bottlenecks) [Topping et al., 2022], and under-reaching (insufficient depth to capture necessary dependencies) [Alon and Yahav, 2021]. Other issues of GNNs come from limited expressivity in distinguishing specific substructures in the graphs, as classic MPNNs are at most powerful as the 1-WL (Weisfeiler-Lehman) test [Xu et al., 2019].
To address these issues, researchers have explored several directions.
Enhancing expressivity. Techniques such as higher-order message passing [Hajij et al., 2023] extend the expressive power beyond the 1-WL test.
Deeper architectures and Graph Transformers (GTs). Alternative methods have been considered, such as Graph Transformers (GTs) [Shi et al., 2021], a class of GNNs that utilize a self-attention mechanism, allowing nodes to interact with any node in the graph, not limited by the underlying graph topology. As a result, GTs do not rely on increased depth to capture long-range dependencies, making it easier to model such relationships. However, this advantage comes with the drawback of quadratic time complexity with respect to the number of nodes, limiting their scalability to larger graphs. Although GTs can be viewed as the solution, Cai et al. [2023] shows that they are not inherently more expressive than MPNNs equipped with a virtual node connected to all others — an alternative that enables global communication in the graph. Then, in practice, GTs do not consistently outperform classic MPNNs, such as GCN [Kipf and Welling, 2017], on the Long-Range Graph Benchmark (LRGB) [Tönshoff et al., 2023], raising doubts about their superiority for long-range propagation tasks. Furthermore, MPNNs and GTs rely heavily on normalization or dropout layers [Luo et al., 2024] to achieve high accuracy, obscuring the model’s internal dynamics.
Graph rewiring and Structure Learning. Methods like spectral rewiring [Gutteridge et al., 2023] and structure learning [Chen et al., 2020] reshape the topology to ease information flow and mitigate over-squashing.
ODE-inspired approaches. Continuous-depth models based on Graph Neural ODEs [Poli et al., 2021] provide an interpretable framework and have shown promising results in mitigating separately the over-smoothing or over-squashing phenomena [Di Giovanni et al., 2023, Gravina et al., 2023].
Despite these advances, most of these methods introduce complexity in the architecture to learn new topology, or as said, GT are quadratic in the number of nodes, scale poorly and do not outperform MPNNs in practice; existing ODE-inspired approaches are cheaper in terms of space or time requirements but they typically consider over-smoothing and over-squashing not in a unique framework. On top of this, choosing an optimal pre-defined depth remains an expensive procedure common to all of these approaches. Our work departs by enabling dynamic, per-node(graph) depth selection without altering the underlying graph topology. In this work, we do not consider the expressivity shortcomings of GNNs related to the 1-WL test, and we leave it as future work.
Asynchronous Message-Passing. A growing body of work attempts to jointly tackle over-smoothing and over-squashing via asynchronous message-passing schemes. This means that the input graph topology differs at every layer from the actual one, in order to allow nodes to send messages following specific paths, reducing the negative effects that over-smoothing and over-squashing could cause.
Finkelshtein et al. [2024] introduces Cooperative GNNs where each node selects interaction types per layer, tailoring aggregation to its local context. We took inspiration from this approach to implement the process of nodes deciding whether to exit rather than how they should interact with neighbors. Despite their efficiency effort and enhanced expressivity, their approach still has higher inference time and memory constraints w.r.t. classic MPNNs such as GCN [Kipf and Welling, 2017], and they also require depth as a hyperparameter. Conversely, we remain efficient w.r.t. classic MPNNs both from a time and space perspective.
Similarly, Errica et al. [2023] learn the depth of the architecture, but not on a per-node basis, and couples it with neighbor message functions during training. These methods enhance expressivity but incur high time and memory overhead w.r.t. classic MPNNs such as GCN. However, we can expect that the required number of layers learned at training time could differ from the one required on unseen instances, for instance, we enable the depth selection both during training and testing.
Compared in general with asynchronous message-passing, our design exploits only the original topology, being a synchronous message-passing approach, and results being more efficient in terms of time and space complexity.
Early-Exit Mechanisms in GNNs. Few methods exist that apply early-exit together with GNNs. Spinelli et al. [2021] implements an exit scheme for nodes to stop updating their representations. However, their method does not explicitly address OST and OSQ, but it attempts to prevent only OST by using the early-exit as a sort of escape. In contrast, we allow the nodes to remain in the network without explicit time constraints or regularizers, as long as they require extracting features, and we simply use OST and OSQ thanks to our design. Other works have explored early exit mechanisms in GNNs, but these are typically limited to node classification and do not investigate theoretically OST and OSQ [Xiao et al., 2021, Han et al., 2024]. Xiao et al. [2021] employs a variational expectation-maximization framework to estimate depth for each node, while we utilize the Gumbel Softmax reparametrization trick. They also consider the OST problem, but they do not design any mechanism that can mitigate it in case of remaining in the forward pass.
Han et al. [2024] based their method on the assumption that less connected nodes should exit earlier, in a data-aware, efficient manner. Our approach is end-to-end differentiable and agnostic to connectivity assumptions that may not generalize to any graph learning task.
Early-Exit Neural Networks (General). Early-exit methods in standard deep learning reduce inference cost by attaching auxiliary classifiers at intermediate layers. Training paradigms include deeply supervised nets [Lee et al., 2014], layer-wise pretraining [Hettinger et al., 2017], and independent exit training [Venkataramani et al., 2015]. At inference, early termination is governed by confidence- or entropy-based thresholds [Wang et al., 2017, Xin et al., 2020, Baccarelli et al., 2020, Xin et al., 2021] or threshold-agnostic strategies [Pomponi et al., 2022, Ju et al., 2021]. These techniques inspire our threshold-free, jointly trainable exit mechanism tailored to graph-structured inputs and obviate the need for hand-tuned confidence thresholds. Despite this, we do not draw inspiration from any specific method, but we recognize the benefit that early-exit can have in graph learning tasks.